### v1

# Calibrating Long-form Generations from Large Language Models

Abstract
To enhance Large Language Models’ (LLMs) reliability, calibration is essential—the model’s assessed confidence scores should align with the actual likelihood of its responses being correct. However, current confidence elicitation methods and calibration metrics typically rely on a binary true/false assessment of response correctness. This approach does not apply to long-form generation, where an answer can be partially correct. Addressing this gap, we introduce a unified calibration framework, in which both the correctness of the LLMs’ responses and their associated confidence levels are treated as distributions across a range of scores. Within this framework, we develop three metrics to precisely evaluate LLM calibration and further propose two confidence elicitation methods based on self-consistency and self-evaluation. Our experiments, which include long-form QA and summarization tasks, demonstrate that larger models don’t necessarily guarantee better calibration, that calibration performance is found to be metric-dependent, and that self-consistency methods excel in factoid datasets. We also find that calibration can be enhanced through techniques such as fine-tuning, integrating relevant source documents, scaling the temperature, and combining self-consistency with self-evaluation. Lastly, we showcase a practical application of our system: selecting and cascading open-source models and ChatGPT to optimize correctness given a limited API budget. This research not only challenges existing notions of LLM calibration but also offers practical methodologies for improving trustworthiness in long-form generation.
affilsepx,
Calibrating Long-form Generations from Large Language Models
Figure 1: A comparison between short-form generation and long-form generation. The correctness of the short-form answer can either be true (1) or false (0), while the correctness of the long-form answer is typically a score between 0 and 1. Both of these scores may vary across evaluators due to subjectivity, hence we conceptualize them as a distribution over $[0,1]$.
Figure 2: Overview of our calibration framework. First, we prompt an LLM to produce an answer to a specific question. Next, we assess the answer’s correctness distribution using an evaluator (GPT-4 or a task-specific metric). Subsequently, we determine the model’s confidence distribution through self-evaluation or self-consistency approaches. Finally, we calculate the calibration score by comparing the correctness and confidence distributions against our predefined metrics.
1 Introduction
Large language models (LLMs) are transforming the landscape of AI, yet their tendency to hallucinate—producing factually incoherent or unrealistic responses—poses a significant hurdle for real-world applications (Ji et al., 2022, Zhang et al., 2023b, Maynez et al., 2020, Guo et al., 2023, Sun et al., 2023, Min et al., 2023). One approach to harnessing their potential lies in calibration: augmenting predictions with confidence scores which are aligned with the likelihood of the predictions being correct. Correctness, however, is often subjective and granular, especially for long-form generation tasks like open-ended QA and summarization (Figure 1). For example, in an ambiguous question, an answer that only considers one of the possible interpretations is neither wrong nor totally correct (Stelmakh et al., 2022).
Current confidence estimation methods, on the other hand, predominantly focus on a binary notion of answer correctness and provide confidence scores for the whole sequence which reflect the statement “I am $x$confident that this answer is completely correct”. Such statements don’t necessarily apply to long-form generation, where correctness varies along a continuum. Furthermore, different annotators can assign varied correctness scores to the same answer Bakker et al. (2022), reflecting the ambiguity and subjectivity inherent in these tasks and their criteria. Given this, it becomes clear that representing an answer’s correctness as a single, definitive score is also limiting.
Addressing these challenges, we introduce a general calibration framework that conceptualizes the correctness of model responses as a distribution across scores between $[0,1]$ (Figure 1). We argue that most long-form generations lie somewhere in between, neither completely incorrect (00) nor perfect ($1$), and that representing the correctness as a distribution better allows capturing the subjectivity of human annotations or the variance in automated procedures for estimating correctness. Consequently, we view the confidence scores associated with the various correctness levels as another distribution over the same range $[0,1]$, corresponding to statements of the form “I am $x$confident that this answer is $y$correct”. Calibration is then defined as the degree of alignment between the predicted confidence distribution and the ground-truth correctness distribution obtained by evaluating the answers.
Methodologically, our framework encompasses three core components (Figure 2): estimating the ground-truth correctness distributions, eliciting confidence distributions from LLMs, and measuring calibration between these distributions. We introduce novel methods for each of these components: estimating correctness via prompting LLMs, eliciting confidence of long-form answers by adapting self-evaluation (Chiang and yi Lee, 2023) and self-consistency (Kadavath et al., 2022), and calibration metrics which measure both the divergence between distributions and their efficacy in selecting high-quality answers.
We leverage our framework to measure calibration for several LLMs on multiple datasets across three long-form QA–including ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022)–and one summarization task, CNNDM (Nallapati et al., 2016). Our results show that stronger LLMs like GPT-3.5 don’t necessarily guarantee better calibration, calibration is metric-dependent, and LLMs exhibit better calibration with self-consistency method on factoid datasets than more open-ended datasets. Furthermore, our analysis reveals effective calibration enhancement strategies for long-form generation tasks. Fine-tuning and temperature scaling boost calibration across all metrics, while integrating relevant source documents and combining self-consistency with self-evaluation yield improvements only in part of metrics. Finally, we demonstrate a practical application of our framework: a cost-effective cascading strategy for using LLMs (Chen et al., 2023a). In this approach, open-source models initially handle queries and, based on their confidence levels—assessed using our system—more advanced API models are engaged as needed. This method ensures cost efficiency while maintaining high performance levels.
In essence, our contributions set a new benchmark for evaluating and improving LLM reliability across text generation tasks, offering:
  1. 1.
    A comprehensive calibration framework applicable to all text generation tasks, enhancing LLM evaluation for critical applications.
  2. 2.
    Innovative methods for confidence elicitation and calibration measurement, applied to a variety of LLMs.
  3. 3.
    Evidence that calibration can be augmented through model fine-tuning, document incorporation, temperature scaling, and mixed elicitation methods.
  4. 4.
    A cost-effective model usage strategy, illustrating the practicality of our calibration system in optimizing LLM deployment.
2 Related Work
Measuring Calibration Calibration (Guo et al., 2017, Minderer et al., 2021), the alignment of a model’s confidence with the actual accuracy of its predictions, has been widely studied in language models, whose probabilities derived from logits are generally found to not be calibrated (Jiang et al., 2020, Kadavath et al., 2022, Chen et al., 2023d). Standard metrics to measure the calibration include Expected Calibration Error (ECE) for confidence-accuracy disparity (Naeini et al., 2015), Brier Score for mean squared prediction-outcome differences, and AUROC for assessing confidence-based correct answer identification (Boyd et al., 2013, Kuhn et al., 2023). Selective Accuracy@Coverage measures accuracy within the model’s most confident predictions (Liang et al., 2023, Cole et al., 2023). However, these metrics are predominantly based on a binary conception of correctness, rendering them less effective for long-form text generation tasks where correctness is better represented as a distribution over a range of scores.
Improving Calibration Several studies develop improved methods for estimating confidence levels in model responses to enhance the calibration of LMs, such as training external calibrators (Mielke et al., 2022), fine-tuning models to express confidence (Lin et al., 2022), and using sensitivity to in-context learning as a proxy for confidence (Chen et al., 2023c). Given the rise of proprietary LLMs with limited access, confidence elicitation methods that don’t require access to logits (Hu et al., 2023) or embeddings have been extensively studied. These methods include verbalization-based, which directly asks the model to output its confidence (Lin et al., 2022), consistency-based, which samples multiple responses and uses consistency as a surrogate for confidence (Kadavath et al., 2022, Kuhn et al., 2023, Cole et al., 2023, Tian et al., 2023a), and the hybrid of both (Xiong et al., 2023, Tian et al., 2023b, Chen and Mueller, 2023). However, most existing methods still operate under the binary assumption of answer correctness. Consequently, confidence is a score that represents the model’s belief that the answer is all correct, inadequate for the complexity of long-form generation tasks.
3 Calibration Formulation
In the following sections, we first formalize the long-form generation calibration problem (§ 3.1), and then we introduce three core components in our calibration framework (Figure 2): the correctness distributions of the answers (§ 3.2), the confidence distributions of the LLM on its answers (§ 4), and the calibration metric to measure the difference between correctness distributions and confidence distributions (§ 5).
3.1 Formulation
Given a dataset $\mathcal{D}$, the model’s answer for each question $Q_{i}$ in the dataset is answer $A_{i}$ (where $i$ indexes the questions in the dataset). To measure how calibrated the model is, we need three steps. First, we apply an evaluator to get the ground-truth correctness distribution $P_{G_{i}}$ where $G_{i}$ is the random variable that denotes the correctness score in answer $A_{i}$. Formally,
$P_{G_{i}}(x)=\text{Pr}(A_{i}\text{ is }x\%\text{ correct})$ (1)
for $x\in[0,1]$. Second, we leverage a confidence elicitation method to extract confidence distribution $P_{C_{i}}$ from an LLM $\mathcal{M}$, where $C_{i}$ is the random variable that denotes the confidence score of the model $\mathcal{M}$ in answer $A_{i}$. We further assume that the confidence scores are normalized across $x$ such that they form a valid distribution, with the same domain as the correctness distribution. Formally, for $x\in[0,1]$
$P_{C_{i}}(x)$ $=\text{Pr}(\mathcal{M}\text{ thinks }A_{i}\text{ is }x\%\text{ correct})$ (2)
Finally, we use $\mathbb{P_{G}}(\mathcal{D})=\{P_{G_{1}},P_{G_{2}},...,P_{G_{n}}\}$ and $\mathbb{P_{C}}(\mathcal{D})=\{P_{C_{1}},P_{C_{2}},...,P_{C_{n}}\}$ to denote the set of correctness and confidence distributions across the whole dataset, where $n$ is the number of data points. In our framework, a model is calibrated if its confidence distribution for its generated response closely matches the ground-truth correctness distribution, as measured by a metric $M(\mathbb{P_{G}}(\mathcal{D}),\mathbb{P_{C}}(\mathcal{D}))$.
3.2 Groundtruth Correctness
A key challenge in the above formulation lies in obtaining a reliable measure of ground-truth correctness $G(A_{i})$. For short-form QA, correctness can be estimated via simple string comparison metrics Rajpurkar et al. (2016), but for long-form answers, the most reliable evaluation is usually via human annotation. For automatic evaluation, traditional metrics like BLEU and ROUGE are known to have limitations for evaluating long-form QA Liu et al. (2023a), and in our preliminary explorations, we found task-specific metrics like the disambiguation-F1 metric for ASQA Stelmakh et al. (2022) to correlate poorly with the quality of responses generated by closed-book LLMs prompted without reference texts.
Hence, in this paper, we adopt the GPT-4 model to evaluate answers, which has been shown to correlate with human judgments in long-form generation (Liu et al., 2023a). We ask GPT-4 to range an answer from 0 (worst) to 5 (best). The evaluation template includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer. See Appendix E.1 for details. To address potential scoring variability caused by the criteria ambiguity and the inherent uncertainty of LLM, we ask the evaluator to repeatedly evaluate each answer, creating a distribution of scores that more accurately represent the answer’s correctness. Specifically, given an LLM evaluator $\mathcal{E}$, $N$ evaluations from it, and a score $x\in[0,1]$:
$P_{G_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{E}(A_{i})_{j}=x)$ (3)
When N set to 1, the correctness distribution becomes a point-mass distribution, typically observed in tasks with objective ground-truth correctness.
4 Confidence Elicitation
We develop two methods to access the model’s confidence distribution about its own answer. The first is to explicitly ask the model to self-evaluate its own answer. The second is to implicitly assess the model’s confidence in the answer by sampling multiple responses from it and estimating the consistency between them. These methods circumvent the need for logit-based approaches (Jiang et al., 2020), which often have limitations. Specifically, the normalization of logits-derived probabilities often leads to a skewed distribution, as most small values cluster at the lower end, reducing the model’s discriminative ability and complicating interpretation. Additionally, logit-based methods generally imply an overestimation of confidence (Mielke et al., 2022). Finally, the inaccessibility of model logits in many commercial, closed-source platforms restricts the practicality of using logit-based methods in these contexts.
4.1 Self Evaluation
We prompt the model to perform self-evaluation in order to obtain its estimated confidence distribution. The self-evaluation template closely mirrors the template for correctness distribution but omits the reference answer (see Appendix E.3 for details). To get the confidence distribution, we ask the model to repeatedly evaluate the same answer with sampling decoding, with the subsequent scores interpreted as the confidence distribution. Formally, given an LLM $\mathcal{M}$, $N$ self-evaluations from it, and a score $x\in[0,1]$:
$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{M}(A_{i})_{j}=x)$ (4)
Such a sampling method provides a more authentic reflection of the model’s internal distribution than logits. By asking the model to assess an answer multiple times, we capture a range of scores that more accurately represent the model’s confidence, which enhances the reliability of the confidence estimation.
4.2 Self Consistency
Another key indicator of model confidence is the consistency among multiple responses a model provides for a given question. Formally, given a primary answer $A_{i}$, other $N$ answers $A_{i}^{1}...A_{i}^{N}$ sampled from LLM $\mathcal{M}$, a metric for measuring the similarity between two answers $Sim(\cdot,\cdot)$, and a score $x\in[0,1]$:
$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(Sim(A_{i},A_{i}^{j})=x)$ (5)
Prior research (Kadavath et al., 2022) indicates that such confidence scores correlate to accuracy when the language model is sufficiently large. While it’s straightforward to determine if two short answers are consistent, evaluating them in long-form generation is more complex. To adapt self-consistency to long-form generation, we introduce four different variations to measure answer similarity in long-form content. 1. Naive: An additional LLM assesses the similarity between two responses. 2. Split: The first answer is divided into individual sentences, and another LLM evaluates the presence of similar statements in the second answer. The similarity score is the average of these sentence comparisons. 3. Claim: Building on the Split method, this variation uses a factual detector to identify factual claims in the first answer and checks for their presence in the second response. 4. Named Entity Recoginition (NER): This method further refines the analysis by comparing the overlap of named entities between two responses. For a detailed description of these methodologies and the models used, refer to the Appendix B.
Each of these methods caters to different scenarios and levels of analysis depth, from the broad-strokes assessment of the naive approach to the detail-oriented NER method. The choice of method depends on the specific requirements of the task and the desired balance between comprehensiveness and specificity in the consistency analysis.
5 Calibration Metric
We propose three distinct calibration metrics to capture how well-calibrated the model is from different perspectives. The first metric Correlation examines the relationship between the expected confidence and correctness scores across the entire dataset, serving as an indicator of the model’s proficiency in ranking answers. The second, Wasserstein Similarity, measures the congruence between the confidence and correctness distributions for each instance in the dataset, offering a nuanced view of the model’s ability to mirror the shape and spread of the actual correctness in its confidence distribution. The third metric Selective F1 assesses the model’s ability to judiciously choose when to respond, reflecting the utility of the confidence scores in picking out the best responses across the dataset. These metrics, detailed in the following sections, provide diverse insights into calibration, enriching our understanding of LLM reliability from multiple angles.
5.1 Correlation
We assess the dataset-level calibration by measuring the correlation between expected values of confidence and correctness distributions. For each answer $A_{i}$ within the dataset, we can calculate the expected confidence score $E(C_{i})$ and the expected correctness score $E(G_{i})$. In the whole dataset, we can get a list of expected confidence scores $\mathbf{E_{C}}$ and a list of expected correctness scores $\mathbf{E_{G}}$. Then we can measure the correlation between them.
$\rho(\mathcal{D})=Corr(\mathbf{E_{C}},\mathbf{E_{G}})$ (6)
where $Corr(\cdot)$ represents the correlation function (Pearson correlation in our experiments). This correlation provides a clear indicator of how well our model’s confidence aligns with its actual correctness across the entire dataset.
5.2 Wasserstein Similarity
Wasserstein distance represents how much “work” it takes to transform one distribution into another. Wasserstein similarity $S_{W}$ adapts the concept of Wasserstein distance $D_{W}$ in a way that it represents similarity rather than distance between two probability distributions. Formally, the instance-level Wasserstein similarity for an answer $A_{i}$ is:
$S_{W}(A_{i})=1-\frac{D_{W}(G_{i},C_{i})}{d_{max}}$ (7)
where $d_{max}$ is the maximum distance in the metric space. And the dataset-level calibration score can be obtained by averaging the instance-level calibration score.
$S_{W}(\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{i}{S_{W}(A_{i})}$ (8)
5.3 Selective F1
In selective answering (e.g., Kamath et al., 2020), models only respond when confident about their accuracy, enhancing their reliability. To assess the effectiveness of LMs in selective answering, we propose the selective F1 metric ($F1_{\tau_{s}}$). This metric quantifies the model’s aptitude in filtering out answers that meet or exceed a predefined correctness threshold, denoted as $\tau_{s}$. Alongside this, we define two critical components: selective precision ($P_{\tau_{s}}$) and selective recall ($R_{\tau_{s}}$). Selective precision calculates the proportion of selected answers that surpass the correctness threshold $\tau_{s}$ relative to the total number of selected answers. In contrast, selective recall compares the number of selected answers meeting this criterion against the total number of correct answers in the dataset that exceed the threshold $\tau_{s}$.
Our approach utilizes a dual-threshold system, consisting of a confidence threshold ($\tau_{c}$) and the correctness threshold ($\tau_{s}$), allowing the model to answer questions only if its confidence in the answer’s expected correctness score exceeding $\tau_{s}$ surpasses $\tau_{c}$. To more formally define the selective F1, let $\mathcal{A}=\{A_{1},A_{2},...,A_{n}\}$ denote the total set of model’s answers in the dataset and $\mathcal{A}^{*}=\{A_{i}\in\mathcal{A}\mid\text{if }\sum_{s\geq\tau_{s}}C_{i}(s)\geq\tau_{c}\}$ denote the set of selected answers. Let the indicator function $I_{\tau_{s}}(A_{i})$ indicate if the expected correctness score $E[G_{i}]$ of $A_{i}$ exceeds $\tau_{s}$
$I_{\tau_{s}}(A_{i})=\begin{cases}1&\text{if }E[G_{i}]\geq\tau_{s}\\ 0&\text{Otherwise}\end{cases}$ (9)
The selective precision on the dataset $\mathcal{D}$ is:
$P_{\tau_{s}}(\mathcal{D})=\frac{\sum_{A_{i}\in\mathcal{A}^{*}}I_{\tau_{s}}(A_{i})}{|\mathcal{A}^{*}|}$ (10)
The selective recall is:
$R_{\tau_{s}}(\mathcal{D})=\frac{\sum_{A_{i}\in\mathcal{A}^{*}}I_{\tau_{s}}(A_{i})}{\sum_{A_{i}\in\mathcal{A}}I_{\tau_{s}}(A_{i})}$ (11)
The selective F1 takes both recall and precision into account:
$F1_{\tau_{s}}(\mathcal{D})=2\frac{P_{\tau_{s}}(\mathcal{D})R_{\tau_{s}}(\mathcal{D})}{P_{\tau_{s}}(\mathcal{D})+R_{\tau_{s}}(\mathcal{D})}$ (12)
ASQA ELI5
Model Method Corr Wass $F1_{0.8}$ Score Corr Wass $F1_{0.8}$ Score
Llama2-13b Self Eval 18.1 71.5 51.0 51.3 21.8 75.0 39.6 53.8
Self Cons 48.0 75.8 3.6 28.8 61.0 39.8
Llama2-70b Self Eval 16.4 64.9 60.3 59.4 18.8 78.9 54.9 61.7
Self Cons 44.3 77.0 29.7 15.5 71.4 51.4
Vicuna-13b Self Eval 30.2 70.8 30.0 50.8 31.2 76.6 36.7 53.0
Self Cons 58.2 72.0 37.6 35.3 63.8 36.9
GPT-3.5-turbo Self Eval 3.6 74.3 83.7 72.6 26.2 76.9 63.5 63.0
Self Cons 30.5 77.7 58.4 18.5 68.6 55.2
QAMPARI CNNDM
Model Method Corr Wass $F1_{0.4}$ Score Corr Wass $F1_{0.8}$ Score
Llama2-13b Self Eval 25.7 66.6 29.0 13.3 19.5 96.5 92.0 77.0
Self Cons 46.0 71.6 42.3 9.6 73.2 24.5
Llama2-70b Self Eval 18.6 41.9 30.1 14.6 6.4 96.9 93.6 77.6
Self Cons 17.1 61.0 39.1 22.7 72.4 20.2
Vicuna-13b Self Eval 16.2 62.6 27.0 11.4 3.9 80.9 64.2 78.2
Self Cons 49.5 78.7 39.7 18.9 78.5 41.7
GPT-3.5-turbo Self Eval 28.7 57.6 53.1 24.0 17.0 96.2 94.7 78.2
Self Cons 44.4 65.8 56.0 18.9 77.8 32.1
Table 1: Calibration Performance Across Four Tasks (in %): “Corr” for Correlation, “Wass” for Wasserstein Similarity, “$F1_{\tau_{s}}$” for Selective F1 Score at threshold $\tau_{s}$, and “Score” representing expected correctness. “Self Eval” for self-evaluation and “Self Cons” for self-consistency. Bold indicates the best score. Key insights include: 1) Higher model capability does not ensure better calibration; 2) Models may exhibit calibration on certain metrics but not others; 3) On factoid datasets (ASQA, QAMPARI), self-consistency outperforms self-evaluation.
6 Experiments and Results
6.1 Setup
Models and Data We measure different sized LLMs’ calibration, including Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), vicuna-13b (Zheng et al., 2023), GPT-3.5-turbo, acorss three long-form QA tasks: ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022), and one summarization task: CNNDM (Nallapati et al., 2016). Details of each dataset can be found in Appendix A.
Correctness Evaluation We apply GPT-4 to evaluate the model-generated answers to approximate the ground-truth correctness distributions for ASQA, ELI5, and CNNDM. In QAMPARI, given that the answer is a list of entities, we evaluate using the F1-5 metric (Gao et al., 2023), calculating the F1 score by the exact match with the gold answer list and defining 100% recall for predictions with at least 5 correct answers. In preliminary experiments, we found that samples from the GPT-4 evaluator had very low variance in terms of the final score assigned to the answers, hence in the following results we only use a single sample, resulting in a point mass distribution for correctness.
Confidence Elicitation We contrast self-evaluation and self-consistency methods for confidence assessment across various tasks, employing task-specific strategies to compare similarities. For ASQA, where answers hinge on factual claims, we utilize a self-consistency approach focused on claims. In the case of QAMPARI, which involves listing entities, we adopt self-consistency-NER. For the ELI5 dataset, a more general self-consistency-naive strategy is applied due to the broad nature of the questions. Lastly, for CNNDM, where summaries are composed of key sentences, we implement a self-consistency-split for detailed analysis. These tailored approaches ensure the most effective evaluation of each dataset’s unique characteristics.
6.2 Main Results
Larger models are not necessarily better calibrated. Table 1 reveals a surprising trend: Vicuna-13b, though performing poorly on ASQA, ELI5 and QAMPARI, has the highest correlation with self-consistency confidence on the same datasets. This might be due to the fact that Vicuna-13b is the only model instruction fine-tuned in a supervised fashion, in contrast to the other models using reinforcement learning human feedback. RL finetuning has been shown to lead to miscalibration in LMs by biasing generations towards highly rewarded behaviors (Kadavath et al., 2022). Scaling the temperature could enhance the calibration of LLMs fine-tuned using RL (see § 6.3). Additionally, Llama-2-13b demonstrates a higher correlation than its larger counterpart, Llama-2-70b, in $3$ out of $4$ datasets with both confidence elicitation methods. However, GPT-3.5-turbo, the strongest model, consistently scores the highest in selective F1 across all datasets. Selective F1 metrics blend performance and calibration and tend to yield higher scores for stronger models.
A model can be calibrated on some dimensions but not on others. Calibration metrics offer complementary perspectives. A model may exhibit low correlation but high Wasserstein similarity, or vice versa. For instance, GPT-3.5-turbo on the ASQA dataset shows low correlation ($3.6$) but high Wasserstein similarity ($74.3$) with self-evaluation confidence. In contrast, Vicuna-13b demonstrates the opposite trend on ASQA with self-consistency confidence (correlation: $58.2$, Wasserstein similarity: $72.0$). This discrepancy arises because correlation and Wasserstein similarity capture different calibration aspects: correlation reflects the model’s ability to appropriately rank good and bad answers, while Wasserstein similarity evaluates how closely a model’s confidence mirrors actual correctness.
Self-Consistency is better than Self-evaluation on factoid datasets. In factoid datasets (ASQA and QAMPARI), self-consistency confidence elicitation outperforms self-evaluation for all models. On QAMPARI, all models show better calibration with self-consistency across all four metrics, while on ASQA, the trend holds for all but selective F1. While self-consistency scores can serve as indicators of answer quality, they are not guaranteed to be the same as ground-truth correctness due to the inability to fully grasp the ground-truth criteria. However, the effectiveness of self-consistency diminishes in more open-ended tasks like ELI5 or the CNNDM summarization dataset. As tasks become more subjective, the power of self-consistency in model calibration begins to wane.
6.3 Improving Calibration
We delve into four strategies aimed at enhancing the calibration score of LLMs: fine-tuning the model to do self-evaluation, retrieving source documents, scaling the temperature, and hybrid confidence elicitation strategies.
Fine-tuning Our study examines enhancing model calibration through fine-tuning, which is conducted on the ASQA dataset. We generate a self-evaluation training dataset using GPT-4 to assess answers from various models to ASQA’s training set questions. This dataset comprises inputs of questions and corresponding model answers, with outputs including a score and an explanation for each answer. The Llama2-13b model is then fine-tuned using LoRA (Hu et al., 2021) on this dataset, aiming to improve its scoring accuracy and explanatory capabilities. We also consider the impact of self-evaluation finetuning on the model’s ability to generate responses, and further experiment with finetuning the model on a combined objective for self-evaluation as well as generating answers from the ASQA training set. For comparison, we also fine-tune the model using generation data only. See Appendix D.1 for details of the training methodology. As Table 2 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task, as well as the limitation of LORA fine-tuning. Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’). The model becomes more confident in terms of self-consistency after fine-tuning.
[…]
Self-Evaluation
None 18.1 $\uparrow$ 71.5 $\uparrow$ 51.0 $\uparrow$ 51.3 $\uparrow$
Evaluation 13.6 $\downarrow$ 67.0 $\downarrow$ 62.0 $\uparrow$ 49.2 $\downarrow$
Generation 20.0 $\uparrow$ 72.4 $\uparrow$ 52.8 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 23.9 $\uparrow$ 74.2 $\uparrow$ 46.6 $\downarrow$ 50.1 $\downarrow$
Self-Consistency
None 48.0 $\uparrow$ 75.8 $\uparrow$ 3.6 $\uparrow$ 51.3 $\uparrow$
Evaluation 46.9 $\downarrow$ 76.7 $\uparrow$ 10.7 $\uparrow$ 49.2 $\downarrow$
Generation 58.9 $\uparrow$ 76.4 $\uparrow$ 32.4 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 54.5 $\uparrow$ 78.4 $\uparrow$ 18.9 $\uparrow$ 50.1 $\downarrow$
Table 2: Comparison among raw and fine-tuned Llama-2-13b on ASQA. “None” for the untrained model, “Evaluation” for the model fine-tuned with the self-evaluation dataset, “Generation” for the model fine-tuned with the ASQA generation data, and “Eval+Gen” for the model fine-tuned with the hybrid dataset combined by self-evaluation dataset and generation data.
Source Documents We examine how additional contextual information influences LLM calibration. Using the Llama-2-13b-chat and GPT-3.5-turbo on ASQA dataset, with a focus on self-consistency as our confidence eliciation method, we test the model’s performance when supplemented with different types of source documents. We experiment with two types of documents: random documents related to the question and ‘oracle’ documents directly relevant to the answers, as included in the dataset release Stelmakh et al. (2022). Our findings, illustrated in Table 3, reveal that oracle documents can enhance model performance and calibration across two out of three metrics for both models, while random documents may be less effective. These results underscore the importance of relevant contextual information in model calibration.
[…]
R 39.0 $\downarrow$ 75.6 $\downarrow$ 30.6 $\uparrow$ 59.6 $\uparrow$
O 49.5 $\uparrow$ 75.6 $\downarrow$ 20.1 $\uparrow$ 65.8 $\uparrow$
[…]
R 36.6 $\uparrow$ 71.8 $\downarrow$ 68.2 $\uparrow$ 64.8 $\downarrow$
O 27.0 $\downarrow$ 79.5 $\uparrow$ 72.8 $\uparrow$ 75.8 $\uparrow$
Table 3: How source documents affect the calibration score. In the document column, “N” means no documents in the input prompt, “R” means randomly selected documents relevant to the topic, and “O” means the oracle documents relevant to the answer. $\uparrow$ denotes that the calibration score goes up when adding documents, while $\downarrow$ means going down.
Temperature We adjust the generation temperature for Llama2-13b from 0.2 to 1 to examine its impact on calibration. The result in Figure 3 reveals a consistent upward trend in all calibration metrics. Notably, the model’s performance initially improves and then deteriorates. This observation implies that modulating the generation temperature can enhance the calibration of the reinforcement-learned model, which is consistent with previous findings observed in short-form answer generation (Kadavath et al., 2022).
[…]
Figure 3: Calibration varies with temperature scaling.
Hybrid Confident Elicitation Prior work (Xiong et al., 2023, Tian et al., 2023b) found that combining self-evaluation and self-consistency can yield a more accurate confidence score. In our study, we explore whether this finding applies to long-form generation. Specifically, for an answer $A_{i}$, we merge confidence distributions obtained through both self-evaluation ($C^{eval}_{i}$) and self-consistency ($C^{consist}_{i}$). The resultant hybrid confidence distribution, $C^{hybrid}_{i}=\alpha C^{eval}_{i}+(1-\alpha)C^{consist}_{i}$, is modulated by the parameter $\alpha$, allowing us to adjust the influence of each distribution. Our experiments are conducted on the ASQA dataset and utilize the correlation as the calibration metric. As shown in Figure 4, we observe that the correlation between confidence and correctness initially increases but then declines as $\alpha$ is varied from 0 to 1. However, this trend is not observed in other metrics (See Figure 7), indicating that while hybrid calibration elicitation may enhance calibration in terms of correlation, it may not have the same impact on other dimensions.
[…]
Figure 4: Correlation vs. Weight Variation for hybrid confidence elicitation. $\alpha$ is the weight for self-evaluation confidence. $\alpha=0$ for solely self-consistency and $\alpha=1$ for solely self-evaluation.
6.4 Application
We showcase a cost-effective application of our system (illustrated in Figure 5) by using a cascade of language models with varying capabilities for selectively responding to queries, aiming for optimal performance within a limited API budget. We initially employ a set of calibrated open-source LMs (selected by our calibration framework) to handle queries. Complex queries, identified by the lower confidence of these models (derived from our self-consistency confidence elicitation method), are escalated to more advanced API LMs. This tiered approach strikes a balance between cost-efficiency and performance in LLM applications.
Figure 5: The illustration of LLM Cascade.
Our experiments, conducted on the ASQA and QAMPARI datasets, involve cascading a series of calibrated open-source models, culminating with GPT-3.5-turbo. See Appendix D.3 for model selection and cascading details. We begin with the first open-source model, addressing questions where it believes the answer has a probability higher than $\tau_{c}$ that the answer’s correctness score is higher than $\tau_{s}$=0.6. Subsequent questions are then processed through the remaining open-source models with the same method. Any queries not confidently handled by these models are ultimately directed to GPT-3.5-turbo for resolution. By adjusting $\tau_{c}$ from 0 to 1, we can vary the number of queries that will be sent to GPT-3.5-turbo, allowing us to analyze the quality of the final answer as the budget increases for calling the API. We use the strongest open-source model’s results as a baseline for a zero API budget, and GPT-3.5-turbo’s results for a full budget scenario. As a baseline, we compare to a strategy where, within a limited number of API requests, a random selection of queries is addressed by GPT-3.5-turbo and the rest are handled by the strongest open-source model. For each API budget, we randomly select questions for GPT-3.5-turbo using 10 different random seeds and calculate the mean and std for the results. As Figure 6 shows, our LLM cascade method always outperforms the random response strategy across various budget levels. We don’t consider the API cost for obtaining self-consistency confidence scores from open-source models, since GPT-3.5 could be substituted with other open-source NLI models for verifying sentence presence in the second answer.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111606060606565656570707070BudgetCorrectnessASQA
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111151515152020202025252525BudgetCorrectnessQAMPARI
Figure 6: Variation in Expected Correctness Score with API Budget Allocation on ASQA and QAMPARI: This graph contrasts the LLM Cascade approach (blue) with the Random Response strategy (red). The red line shows the mean correctness score from 10 random seeds, while the light red shaded area indicates the range of standard deviation. Notably, a budget of 1 corresponds to the correctness achieved by GPT-3.5-turbo, while a budget of 0 reflects the performance of the best-performing open-source model selected for LLM cascade.
7 Conclusion
In conclusion, our research introduces a comprehensive calibration system designed to evaluate LLMs across both short-form and long-form text generation. We have developed confidence elicitation methods that derive confidence distributions from LLMs. To assess calibration, we introduced three new metrics, each tailored to capture different aspects of LLM performance. Our findings reveal that larger LLMs do not necessarily equate to better calibration, and the degree of calibration significantly varies across different datasets. This challenges prevailing assumptions in the field and underscores the need for nuanced approaches in LLM development and assessment. Furthermore, our study explores various strategies to enhance LLM calibration, offering practical solutions to common issues faced in the field. We also present a potential application of our system, demonstrating its effectiveness in optimizing LLMs performance within the constraints of limited API budgets. The system and tools we present are crucial for further improving the liability of LLMs.
Our research highlights a couple of research directions, including (1) Advancing Calibration Techniques. Develop more sophisticated methods like integrating human feedback into the calibration process, incorporating machine learning algorithms that can adaptively adjust calibration based on the context or type of query, and so on. (2) Cross-Domain Applications. Extend the application of our system to other domains such as medical text generation, legal document analysis, education chatbots and so on. (3) Real-time Calibration. Develop methods for real-time calibration of LLMs in dynamic environments (e.g. interactive chatbots, updated knowledge base). (4) Pre-training with Calibration. Incorporate calibration into the pre-training phase of LLMs. (5) Ethical and Biases Consideration. Investigating how different calibration strategies affect the fairness and bias in LLM outputs can be crucial for responsible AI. (6) Scalability and Efficiency. Explore ways to improve the scalability and efficiency of our calibration system, making it applicable to large-scale, real-world scenarios.
Appendix A Dataset
ASQA (Answer Summaries for Questions which are Ambiguous) (Stelmakh et al., 2022) is a specialized long-form factoid dataset, designed to address ambiguous factoid questions that yield different correct answers based on their interpretations. This dataset challenges models to synthesize factual information from multiple sources, creating coherent long-form summaries that effectively resolve the inherent ambiguities in these questions.
ELI5 (Fan et al., 2019) is a comprehensive open-ended long-form dataset, encompassing over 270,000 threads from the Reddit forum “Explain Like I’m Five.” This unique platform features community-generated responses to a wide array of questions, all tailored to be easily understandable by a five-year-old audience. The majority of queries in ELI5 are centered around ’how,’ ’why,’ and ’what’ questions, which necessitate comprehensive, detailed responses supported by evidence from multiple passages.
QAMPARI (Amouyal et al., 2022) is a factoid dataset where answers are presented as lists of entities dispersed across multiple paragraphs. Its construction involves an automated process that utilizes Wikipedia knowledge graphs and tables. Questions are manually paraphrased, and answers are thoroughly verified for accuracy. Notably, each question in QAMPARI is associated with an average of 13 answers, demonstrating its breadth.
CNNDM (Nallapati et al., 2016) is a large-scale news summarization dataset containing news articles from CNN [^1] https://www.cnn.com/ and DailyMail [^2] https://www.dailymail.co.uk/. The original CNNDM dataset consists of both source news articles and reference summaries. However, recent work Liu et al. (2023b); Zhang et al. (2023a) has found that the provided reference summaries are not of very good quality and zero-shot LLMs summaries are preferred by human annotators over the reference summaries.
In the ACLE (Automatic LLMs’ Citation Evaluation) (Gao et al., 2023), a pioneering benchmark for assessing LLMs’ citation capabilities, a subset of 1,000 examples is randomly selected from the development sets of ASQA, ELI5, and QAMPARI to form a test set for each task. For our specific analysis, we choose to utilize the first 500 examples from each of these datasets in ACLE as our test set, providing a focused and representative sample for each task. For CNNDM, we utilize 100 examples as our test set.
Appendix B Self-consistency
We propose four different self-consistency based methods tailored for long-form generation, each with a different strategy to measure the similarity between two long-form answer.
Naive The most basic approach utilizes an additional LLM (GPT-3.5-turbo in our experiments, which can be replaced by other models trained for this task) to determine if two responses are akin, assigning a corresponding similarity score. This method diverges from the relevant technique in contemporary research (Chen et al., 2023b), which primarily focuses on identifying the most consistent answer. Instead, our approach aims to secure specific consistency ratings that reflect the model’s assurance in its primary answer, offering a general overview of the answers’ similarity.
Sentence Split For a more detailed similarity analysis between the two answers, we split the first answer into individual sentences. Another LLM (GPT-3.5-turbo or a similar NLI model) is then used to evaluate whether similar statements are present in the second answer. This method’s limitation is that not all sentences carry equal informational weight. Some may be filler or less informative, potentially skewing the similarity assessment.
Claim To further refine the approach, we focus on sentences that make factual claims. This involves two steps: first, using a claim detector to identify factual claims within a sentence and then using an NLI model to determine if similar factual claims exist in the second answer. This method operates under the assumption that factual claims are the most critical components of an answer, representing its core information. We leverage a DeBERTa-V2 (He et al., 2020) trained by ClaimBuster as our fact detector and GPT-3.5 as the NLI models.
Named Entity Recognition Advancing the granularity further, we compare named entities between two responses. We identify and compare entities present in both answers by utilizing a named entity recognition model. The degree of overlap in these entities serves as an indicator of answer similarity. This approach focuses more on concrete, identifiable elements within the answers. We use a Roberta-large (Liu et al., 2019) trained with SpanMarker framework  [^3] https://github.com/tomaarsen/SpanMarkerNER, which can be replaced by other NER models.
Appendix C Main Experiment
We assess the calibration of variously sized models (Llama-2-13b, Llama-2-70b, Vicuna-13b, GPT-3.5-turbo) across three long-form QA datasets (ASQA, ELI5, QAMPARI) and one summarization dataset (CNNDM). The process involves several steps: 1. Generation: Asking the model to generate answers for questions in the dataset 2. Correctness Assessment: We utilize GPT-4 to evaluate the correctness of the models’ answers, except for QAMPARI, where we directly apply the F1-5 metric. 3. Confidence Distribution Derivation: After generating answers, models employ self-evaluation or self-consistency methods to derive their confidence distribution. 4. Calibration Measurement: .
C.1 Generation
In the answer generation phase, we employ a 3-shot in-context learning approach for each long-form QA dataset, providing three exemplars to guide the models. For the CNNDM dataset, we adopt a 0-shot strategy, aligning with prior studies (Goyal et al., 2022). The generation of answers utilizes a top-K sampling method, setting the generation temperature for all models at 0.6 and the top-K parameter at 10.
C.2 Correctness Evaluation
For ASQA, ELI5, CNNDM, we ask the GPT-4 to evaluate the correctness of examples. For ASQA and ELI5, We ask GPT-4 to evaluate the answer three times, producing a distribution capturing criteria ambiguity and model’s inherent subjectivity. However, for CNNDM, we only ask GPT-4 to evaluate once to save computation given that most answers correctness concentrate around 0.8 with small variance. The GPT-4 evaluation template can be found in Appendix E.1 and criteria for different tasks can be found in Appendix E.2. For QAMPARI, which involves generating a list of entities as answers, we determine that F1 scores provide a more suitable measure of correctness than GPT-4 evaluations. To this end, we adopt the F1-5 metric (Gao et al., 2023), computing the F1 score based on direct matches with the correct answer list and assigning a recall of 100% for responses containing at least five accurate answers (recall-5).
C.3 Confidence Elicitation
After getting the answer from the model, we leverage the self-evaluation or self-consistency method to derive confidence distribution from the model.
Self-Evaluation We prompt the model to self-evaluate an answer 10 times, creating a confidence distribution based on its self-evaluation scores. Each self-evaluation includes task instructions, grading criteria, and evaluation examples (three for all long-form QA datasets and one for CNNDM due to length constraints). Additionally, it contains specific instructions for self-evaluation, incorporating both the question and the answer under evaluation. The detailed self-evaluation template and criteria are available in Appendix E.3 and Appendix E.2 respectively.
Self-consistency In the self-consistency approach, we generate an answer to the same question 10 times, designating the first response as the primary answer. We then calculate similarity scores between the primary answer and the remaining responses. These scores collectively create the model’s confidence distribution for the primary answer. For various datasets, tailored strategies are employed to compare similarities between two answers.
For ASQA, we employ a self-consistency-claim approach. To assess the similarity between the two answers, we first identify factual claims in the first answer using a ClaimBuster-trained DeBERTa V2 (He et al., 2020) claim detector. We then verify the presence of these claims in the second answer using GPT-3.5-turbo. The similarity score is the average presence of factual claims across answers.
In QAMPARI, self-consistency-NER is used to determine confidence distribution. As the answers are entity lists, we extract entities by separating commas. The similarity score is calculated based on the proportion of overlapping entities between two answers, relative to the total entities in the first answer.
For ELI5, where answers provide easily understandable explanations, we focus on the overall content. Here, we apply a self-consistency-naive method, asking an additional model (GPT-3.5-turbo) to assign a similarity score to each answer pair.
In CNNDM, where answers are summaries highlighting key points of an article, we gauge the similarity between two answers by evaluating the overlap of key points. To achieve this, we implement the self-consistency-split method. We dissect the first answer into individual sentences and then use GPT-3.5-turbo to determine if each sentence is present in the second answer. The similarity score is derived by averaging the presence of these segmented sentences in the comparative answer.
Appendix D Improving Calibration
D.1 Fine-tuning
In our study, we focus on improving model calibration on the ASQA dataset. We generate a self-evaluation training dataset with GPT-4, which evaluates different models responses to questions drawn from ASQA training set. We create the self-evaluation data in two steps. The initial phase involves the creation of a diverse answer pool. This is achieved by employing a suite of models with varying computational capacities, including Llama-2-7b, Llama-2-13b, Llama2-70b, Vicuna-13b, and ChatGPT. Each model generates responses to a spectrum of questions drawn from the training set of ASQA task, ensuring the resultant answer pool encompasses a broad quality spectrum, from low to high. Subsequently, we employ GPT-4 to critically assess these answers, assigning a score and providing a corresponding explanation for its evaluation. This process yields a rich dataset, each instance of which encompasses a question, a model-generated answer, an evaluative score, and a justification for that scoring. This approach results in a comprehensive dataset with 2,000 self-evaluation examples (1,800 for training and 200 for validation), each including a question, model-generated answer, score, and justification. For generation data, we use 80% of ASQA’s dataset (4,353 examples) for training and the remaining 20% for validation. The hybrid dataset combines the self-evaluation and generation training sets, using the self-evaluation validation set for assessment.
Regarding training, we fine-tune Llama2-13b-chat model using LORA on this dataset. We maintain consistent parameters across all scenarios: a learning rate of 5e-6, five epochs, 100 warm-up steps, and a total batch size of 4 (achieved through 4 gradient accumulation steps across four GPUs, with a batch size of 1 per device).
D.2 Hybrid Confidence Elicitation
We integrate self-evaluation and self-consistency methods to assess their combined impact on model calibration. Experiments are performed on the ASQA dataset using Llama-2-13b, Llama-2-70b, and Vicuna-13b models, adjusting the self-consistency weight from 0 to 1. Our findings indicate that this hybrid approach improves correlation, though it does not significantly affect other metrics.
[…]
Figure 7: Hybrid confidence elicitation
D.3 LLM Cascade
Our research utilizes ASQA and QAMPARI datasets, applying self-consistency confidence in our methodology. We use the initial 500 samples from the evaluation data provided in Gao et al. (2023) as our test set and samples from 700 to 800 as the validation set. Initially, we assess the calibration of models Llama-2-13b, Llama-2-70b, and Llama-2-70b on the validation set, with the findings presented in Table 4. Notably, the Llama-2-70b model exhibits inferior calibration compared to the other models when tested on QAMPARI, leading us to exclude it from the model cascade for the test set. In the cascading process, we prioritize models based on their performance, ensuring that the most capable model addresses questions that multiple models confidently identify. Consequently, the sequence of cascaded models for ASQA is Llama-2-70b, Llama-2-13b, Vicuna-13b, GPT-3.5-turbo, and for QAMPARI, it is Llama-2-13b, Vicuna-13b, GPT-3.5-turbo. Then we perform our pipeline on the test sets for these tasks.
[…]
ASQA
Vicuna-13b 50.3 70.7 49.2 47.4
Llama-2-13b 51.7 76.4 21.3 47.6
Llama-2-70b 48.0 76.0 45.2 60.4
QAMPARI
Vicuna-13b 48.8 75.5 27.3 13.5
Llama-2-13b 43.0 66.2 45.7 14.1
Llama-2-70b 3.4 56.0 20.0 8.9
Table 4: Open-source models results on validation sets of ASQA and QAMPARI
Appendix E Prompts
This section introduces the prompts used for our experiments.
E.1 Correctness Evaluation Template
Similar to contemporary work (Kim et al., 2023), our evaluation template for GPT-4 evaluation to get the ground-truth correctness distribution of an answer includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
{task instruction} You will be given a question, a reference answer, and a student’s answer. Please evaluate the student’s answer based on both your knowledge and the reference answer, and provide a score from 0-5 to the student’s answer. Keep in mind that the reference answer is not the sole correct response. Assess for both factual accuracy and relevance to the question. The following are the scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Reference answer: {reference answer} Student’s answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation.
E.2 Criteria
Below are the criteria for various tasks, with a special note that the CNNDM summarization task utilizes a distinct evaluation template.
ASQA Criterion
5 - Completely Correct and Highly Relevant: The answer fully addresses the question, resolves the ambiguity, and provides a well-rounded resolution. All facts presented in the answer are accurate and relevant. 4 - Mostly Correct and Relevant: The answer is very relevant and addresses the ambiguity well, but might have a minor oversight or inaccuracy. All the facts presented are accurate and relevant, or with only minor errors. 3 - Partially Correct and Relevant: The answer is generally on topic and attempts to address the ambiguity, but there might be inaccuracies or omissions. The majority of the facts are correct, with a few errors. 2 - Flawed but Somewhat Relevant: The answer somewhat addresses the topic but does not fully explore the question’s ambiguity or does not provide a complete resolution. The facts presented are a mix of correct and incorrect information, with about half being accurate. 1 - Mostly Incorrect or Mostly Irrelevant: The answer slightly touches upon the topic but misses the main point. The majority of the facts presented are incorrect, with only a small portion being accurate. 0 - Completely Incorrect or Completely Irrelevant: The student’s answer is completely off-topic, not related to the question at all, or contains only incorrect information.
ELI5 Criterion
5 - Perfectly Addressed, Accurate and Clarity: The answer flawlessly addresses the question with exceptional accuracy and clarity. It simplifies complex concepts effectively and does so in a way that is captivating and memorable. 4 - Accurate and clear: The answer is accurate, relevant to the question, and presented in a way that is engaging and understandable. It simplifies complex concepts effectively but may miss a small opportunity for further clarification or engagement. 3 - Moderately Accurate and Understandable: The answer is mostly accurate and somewhat understandable. It addresses the question reasonably well but may lack detail or contain some inaccuracies. It may use complex terms or concepts that are not broken down into simpler ideas. 2 - Relevant but Lacks Clarity or Accuracy: The answer is related to the question but lacks clarity or contains partial inaccuracies. It attempts to simplify the idea but does not do so effectively, leaving room for confusion or misunderstanding. 1 - Significantly Flawed: The answer addresses the question to a minimal extent but contains significant inaccuracies or misleading information. It might show a basic attempt to simplify the concept but fail in accuracy or relevance. 0 - Completely Inaccurate or Irrelevant: The answer is entirely off-topic, irrelevant, or factually incorrect. It fails to address the question and does not simplify complex ideas.
* Estimated Accuracy Assessment (0-3 Points) * 3 Points: All answers provided seem correct based on available knowledge. * 2 Points: Majority of the answers seem correct. * 1 Point: Some answers are correct, but there are notable inaccuracies. * 0 Points: No answers seem correct or very high degree of inaccuracy. * Estimated Completeness Assessment (0-2 Points) * 2 Points: The response seems comprehensive, covering a broad range of known or expected correct answers. * 1 Point: The response covers some correct answers but misses significant known or expected answers. * 0 Points: The response is highly incomplete, missing most of the known or expected correct answers. * Total Score (0-5 Scale) * Add the points from Estimated Accuracy and Estimated Completeness.
You will be given one summary written for a news article. Your task is to rate the overall quality of the summary with a score from 0 to 5, where 0 is the lowest and 5 is the highest. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Steps: 1. Read the news article carefully and identify the main topic and key points. 2. Read the summary and compare it to the news article. Check if the summary covers the main topic and key points of the news article, and if it presents them in a clear and logical order. 3. Assign a score for the summary quality on a scale of 0 to 5, where 0 is the lowest and 5 is the highest. Task Input: Article: {article} Summary: {summary} Now please provide your score of the summary in the format of “Score: <Your score>/5” and give your explanation.
E.3 Self-Evaluation Template
{task instruction} You will be given a question and a student’s answer. Please evaluate the student’s answer, and provide a score from 0-5 to the student’s answer based on the following scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation. Assess for both factual accuracy and relevance to the question.
E.4 Self-Consistency Template
In self-consistency approaches, assessing the similarity between two answers requires the use of an additional LLM in naive and split/claim methods. For the self-consistency-naive method, this involves an LLM directly providing similarity scores for the pair of answers. Meanwhile, for the self-consistency-split/claim method, another LLM is tasked with detecting the presence of sentences from answer 1 in answer 2. The templates for both the naive method and the split/claim method are outlined below.
You will be presented with a question followed by two answers. Evaluate how similar these answers are, considering their amount of information provided, factual content, effectiveness in addressing the question, format and organization. Conclude by providing an integer similarity score between 0 and 5. Question: {question} Answer 1: {answer1} Answer 2: {answer2} Now please give your similarity score in the format of “Similarity score: <Your score>/5” and give your explanation. Make your answer short and concise.
You will receive a sentence and a response; please ascertain if a similar statement is present in the response. Sentence: {sentence} Response: {response} Verify if the response contains a statement resembling the target sentence. Answer yes or no.
Appendix F Qualitative Analysis
We showcase some successful and failed calibration examples from LLama-2-13b on Qampari and ASQA with self-consistency method.
type content
QAMPARI Successful Example 1
Question
Who directed a film that had P. Balachandran as a screenwriter?
Answer
Kamal, P. Balachandran, T. K. Rajeev Kumar, V. K. Prakash, Rajeev Ravi.
Model Answer
Adu, Ammayum Muyalum, Bhoothakkannadi, Chattambi, Kakkakku Pinnaka…
Other Answer 1
Agraharathile Kazhuthai, Kaliyugam, Kodungallur, Kodiyettam …
Other Answer 2
Kodiyettam, Kuttyedam, Oru CBI Diary Kurippu…
Other Answer 3
Aravindan, K.G. George, John Abraham, Shyam Benegal …
Self-consistency scores
[0, 0, 0, …] (%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $0\%$ Distribution: $\{0:1.0\}$
QAMPARI Successful Example 2
Question
Which cities or municipalities are located in the province of Basilan?
Answer
Akbar, Al-Barka, Hadji Muhtamad, Isabela City, Lamitan, Lantawan, Maluso, Sumisip, Tabuan-Lasa, Tipo-Tipo, Tuburan, Ungkaya Pukan.
Model Answer
Al-Barka, Basilan, Isabela, Lamitan, Tipo-Tipo, Maluso, Sumisip.
Other Answer 1
Basilan, Isabela, Lamitan, Lantawan, Tipo-Tipo, Sumisip.
Other Answer 2
Basilan, Tipo-Tipo, Isabela City, Maluso, Sumisip, Tuburan.
Other Answer 3
Basilan, Isabela, Tipo-Tipo, Lamitan, Sumisip, Al-Barka, Maluso, Tuburan.
Self-consistency scores
[71.4, 57.1, 85.7, …](%)
Correctness
Score: $83.3\%$ Distribution: $\{0.8:1.0\}$
Confidence
Score: $76.2\%$ Distribution: $\{0.6:0.11,0.8:0.78,1.0:0.11\}$
QAMPARI Failed Example 1 (Under Confident)
Question
What city or municipality located in the Philipines are in the Tarlac province?
Answer
Anao, Bamban, Camiling, Capas, Concepcion, Gerona, La Paz, Mayantoc, Moncada, Paniqui, Pura, Ramos, San Clemente, San Jose, San Manuel, Santa Ignacia, Tarlac City, Victoria.
Model Answer
Capas, Concepcion, Gerona, La Paz, Mayantoc, Paniqui, San Jose, Santa Ignacia, Tarlac City, Victoria.
Other Answer 1
Capas, Concepcion, Camiling, Tarlac City, Paniqui, San Jose, Santa Lucia, Victoria.
Other Answer 2
Capas, Tarlac, Concepcion, Tarlac, Gerona, Tarlac, La Paz, Tarlac, Paniqui, Tarlac, San Jose, Tarlac, Tarlac City.
Other Answer 3
Capas, Concepcion, Camiling, Tarlac City, San Jose.
Self-consistency scores
[60, 60, 30, …](%)
Correctness
Score: $100\%$ Distribution: $\{1:1.0\}$
Confidence
Score: $51.1\%$ Distribution: $\{0.4:0.56,0.6:0.33,0.8:0.11\}$
QAMPARI Failed Example 1 (Over Confident)
Question
What are some notables mailing list softwares ?
Answer
Dada Mail, Discourse, GNU Mailman, GroupServer, LISTSERV, Majordomo, Organic Groups Mailing List, Sympa.
Model Answer
Mailchimp, Constant Contact, Campaign Monitor, AWeber, GetResponse, MailPoet, Sendinblue, MailerLite, Klaviyo, ConvertKit.
Other Answer 1
Mailchimp, AWeber, GetResponse, Campaign Monitor, Constant Contact, Klaviyo, ConvertKit, MailPoet, Mailster, Sendinblue, MailerLite.
Other Answer 2
Mailchimp, Constant Contact, AWeber, Campaign Monitor, GetResponse, Klaviyo, MailPoet, Mailster, Sendinblue, MailerLite, ConvertKit, ActiveCampaign, and Constant Contact.
Other Answer 3
Mailchimp, AWeber, GetResponse, Campaign Monitor, MailerLite, Constant Contact, Klaviyo, ConvertKit, Sendinblue, and Emma.
Self-consistency scores
[100, 100, 90, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $97.8\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
Table 5: Examples for QAMPARI
type content
ASQA Successful Example 1
Question
What are the three types of enzymes in digestion?
Answer
Digestive enzymes are a group of enzymes that break down polymeric… Lipases split fatty acids… Proteases and peptidases split proteins … Amylases split carbohydrates … Nucleases split nucleic acids into nucleotides…
Model Answer
There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. Proteases are produced by the stomach and pancreas, while amylases are … These enzymes work in concert …
Other Answer 1
Digestive enzymes are proteins that help … There are three main types of digestive enzymes: proteases, which break down proteins into amino acids; amylases, which break down carbohydrates into simple sugars; and lipases, which break down fats into glycerol and fatty acids. These enzymes are produced by different organs and …
Other Answer 2
There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. These enzymes are produced by the salivary glands, stomach, and pancreas, and work in various parts of the digestive system…
Self-consistency scores
[100, 100, 0, …] (%)
Correctness
Score: $100\%$ Distribution: $\{1.0:1.0\}$
Confidence score
Score: $95.6\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
ASQA Successful Example 2
Question
When did ben rectors new album come out?
Answer
Benjamin Evans Rector is an American singer… He has released seven studio albums, including Magic on June 22, 2018, which peaked at No. 1 .. Brand New which released on August 28, 2015… Older albums such as Something Like This, released on September 13, 2011 and The Walking in Between, came out August 20, 2013.
Model Answer
Ben Rector’s new album, "Magic", was released on March 25, 2022.
Other Answer 1
Ben Rector’s new album, "Brand New", was released on March 10th, 2022.
Other Answer 2
Ben Rector, an American singer-songwriter, released his new album "Magic" on October 20, 2022.
Self-consistency scores
[0, 0, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence
Score: $0\%$ Distribution: $\{0:1.0\}$
ASQA Failed Example 1 (Under Confident)
Question
Who plays charlie on it’s always sunny?
Answer
Charlie Day plays the fictional character Charlie Kelly… Day is an American actor, screenwriter…
Model Answer
Charlie Day plays Charlie on It’s Always Sunny in Philadelphia. He is also known for his role as Dale Arbus … Day is a member of the improvisational comedy …
Other Answer 1
Dennis is played by Glenn Howerton.
Other Answer 2
Charlie on It’s Always Sunny in Philadelphia is played by Charlie Day.
Self-consistency scores
[0, 33.3, …](%)
Correctness
Score: $80\%$ Distribution: $\{0.8:1.0\}$
Confidence
Score: $31.1\%$ Distribution: $\{0:0.22,0.4:0.78\}$
ASQA Failed Example 2 (Over Confident)
Question
Who sings wake me up when it’s all over?
Answer
“Wake Me Up” is a song by Swedish DJ and record producer Avicii. Aloe Blacc, which is the stage name for Egbert Nathaniel Dawkins III provides vocals for the song while Mike Einziger provides acoustic guitar. “Wake Me When It’s Over” is a song written and recorded by country singer Willie Nelson in 1962.
Model Answer
“Wake Me Up When September Ends” is a song written by Billy
Other Answer 1
“Wake Me Up When September Ends” is a song by American rock band Green Day, written by lead vocalist Billie Joe Armstrong…
Other Answer 2
“Wake Me Up When September Ends” is a song by Green Day, written by the band’s lead vocalist and guitarist Billie Joe Armstrong…
Self-consistency scores
[100, 100, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $77.8\%$ Distribution: $\{0:0.22,1.0:0.78\}$
Table 6: Examples for ASQA

### v2

@@
 Abstract
-To enhance Large Language Models’ (LLMs) reliability, calibration is essential—the model’s assessed confidence scores should align with the actual likelihood of its responses being correct.
-However, current confidence elicitation methods and calibration metrics typically rely on a binary true/false assessment of response correctness.
-This approach does not apply to long-form generation, where an answer can be partially correct.
+To enhance Large Language Models’ (LLMs) reliability, calibration is essential—the model’s confidence scores should align with the likelihood of its responses being correct.
+However, traditional calibration methods typically rely on a binary true/false assessment of response correctness, unsuitable for long-form generations where an answer can be partially correct.
 Addressing this gap, we introduce a unified calibration framework, in which both the correctness of the LLMs’ responses and their associated confidence levels are treated as distributions across a range of scores.
-Within this framework, we develop three metrics to precisely evaluate LLM calibration and further propose two confidence elicitation methods based on self-consistency and self-evaluation.
-Our experiments, which include long-form QA and summarization tasks, demonstrate that larger models don’t necessarily guarantee better calibration, that calibration performance is found to be metric-dependent, and that self-consistency methods excel in factoid datasets.
-We also find that calibration can be enhanced through techniques such as fine-tuning, integrating relevant source documents, scaling the temperature, and combining self-consistency with self-evaluation.
-Lastly, we showcase a practical application of our system: selecting and cascading open-source models and ChatGPT to optimize correctness given a limited API budget.
-This research not only challenges existing notions of LLM calibration but also offers practical methodologies for improving trustworthiness in long-form generation.
+We develop three metrics for assessing LLM calibration and propose confidence elicitation methods based on self-consistency and self-evaluation.
+Our experiments demonstrate that larger models don’t necessarily guarantee better calibration, that various calibration metrics complement each other, and that self-consistency methods excel in factoid datasets.
+We also find that calibration can be enhanced through techniques such as fine-tuning, scaling the temperature.
+Finally, we illustrate one application of long-form calibration through selective answering in long-form responses, optimizing correctness within a constrained API budget.
 affilsepx,
@@
 Figure 2: Overview of our calibration framework.
-First, we prompt an LLM to produce an answer to a specific question.
-Next, we assess the answer’s correctness distribution using an evaluator (GPT-4 or a task-specific metric).
-Subsequently, we determine the model’s confidence distribution through self-evaluation or self-consistency approaches.
+We prompt an LLM to produce an answer to a specific question, assess the answer’s correctness distribution using an evaluator (task-specific metric/GPT-4 metric/human metric), and determine the model’s confidence distribution through self-evaluation or self-consistency approaches.
 Finally, we calculate the calibration score by comparing the correctness and confidence distributions against our predefined metrics.
 1 Introduction
-Large language models (LLMs) are transforming the landscape of AI, yet their tendency to hallucinate—producing factually incoherent or unrealistic responses—poses a significant hurdle for real-world applications (Ji et al., 2022, Zhang et al., 2023b, Maynez et al., 2020, Guo et al., 2023, Sun et al., 2023, Min et al., 2023).
-One approach to harnessing their potential lies in calibration: augmenting predictions with confidence scores which are aligned with the likelihood of the predictions being correct.
-Correctness, however, is often subjective and granular, especially for long-form generation tasks like open-ended QA and summarization (Figure 1).
-For example, in an ambiguous question, an answer that only considers one of the possible interpretations is neither wrong nor totally correct (Stelmakh et al., 2022).
-Current confidence estimation methods, on the other hand, predominantly focus on a binary notion of answer correctness and provide confidence scores for the whole sequence which reflect the statement “I am $x$confident that this answer is completely correct”.
-Such statements don’t necessarily apply to long-form generation, where correctness varies along a continuum.
-Furthermore, different annotators can assign varied correctness scores to the same answer Bakker et al.
-(2022), reflecting the ambiguity and subjectivity inherent in these tasks and their criteria.
-Given this, it becomes clear that representing an answer’s correctness as a single, definitive score is also limiting.
-Addressing these challenges, we introduce a general calibration framework that conceptualizes the correctness of model responses as a distribution across scores between $[0,1]$ (Figure 1).
-We argue that most long-form generations lie somewhere in between, neither completely incorrect (00) nor perfect ($1$), and that representing the correctness as a distribution better allows capturing the subjectivity of human annotations or the variance in automated procedures for estimating correctness.
-Consequently, we view the confidence scores associated with the various correctness levels as another distribution over the same range $[0,1]$, corresponding to statements of the form “I am $x$confident that this answer is $y$correct”.
-Calibration is then defined as the degree of alignment between the predicted confidence distribution and the ground-truth correctness distribution obtained by evaluating the answers.
-Methodologically, our framework encompasses three core components (Figure 2): estimating the ground-truth correctness distributions, eliciting confidence distributions from LLMs, and measuring calibration between these distributions.
-We introduce novel methods for each of these components: estimating correctness via prompting LLMs, eliciting confidence of long-form answers by adapting self-evaluation (Chiang and yi Lee, 2023) and self-consistency (Kadavath et al., 2022), and calibration metrics which measure both the divergence between distributions and their efficacy in selecting high-quality answers.
-We leverage our framework to measure calibration for several LLMs on multiple datasets across three long-form QA–including ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022)–and one summarization task, CNNDM (Nallapati et al., 2016).
-Our results show that stronger LLMs like GPT-3.5 don’t necessarily guarantee better calibration, calibration is metric-dependent, and LLMs exhibit better calibration with self-consistency method on factoid datasets than more open-ended datasets.
-Furthermore, our analysis reveals effective calibration enhancement strategies for long-form generation tasks.
-Fine-tuning and temperature scaling boost calibration across all metrics, while integrating relevant source documents and combining self-consistency with self-evaluation yield improvements only in part of metrics.
-Finally, we demonstrate a practical application of our framework: a cost-effective cascading strategy for using LLMs (Chen et al., 2023a).
-In this approach, open-source models initially handle queries and, based on their confidence levels—assessed using our system—more advanced API models are engaged as needed.
-This method ensures cost efficiency while maintaining high performance levels.
-In essence, our contributions set a new benchmark for evaluating and improving LLM reliability across text generation tasks, offering:
-  1. 1.
-    A comprehensive calibration framework applicable to all text generation tasks, enhancing LLM evaluation for critical applications.
-  2. 2.
+Confidence calibration in large language models (LLMs) aims to align the model’s internal confidence with a probabilistic perspective of its answers’ correctness (i.e.
+quality), enhancing reliability and interpretability for aiding human decision-making Kadavath et al.
+(2022).
+People intuitively understand and utilize probabilities Cosmides and Tooby (1996), making this approach crucial for practical applications.
+Conventional calibration Guo et al.
+(2017) treats answer correctness as binary (true or false) and seeks to align the model’s confidence with the likelihood of model’s answer being correct, typically stated as: “I am xconfident that this answer is completely correct.” However, the correctness of long-form generation is not always either true or false but can be partially correct (Figure 1).
+Therefore, a single confidence score for long-form outputs is ambiguous: it can either imply “I am xconfident that the answer is 100% correct” or “I am 100% confident that the answer is xcorrect.” The former fails to capture the graded notion of long-form answer correctness, while the latter focuses on self-evaluation of correctness, rather than calibration as it overlooks the confidence at specific correctness levels.
+Addressing this challenge, we propose to conceptualize the model’s confidence as distribution across scores between $[0,1]$ to capture the nuanced understanding of the model of each correctness level of the long-form answer, corresponding to the statement “I am $x$confident that this answer is $y$correct”.
+Moreover, we also view the correctness of an LLM’s response as a distribution across scores between $[0,1]$ to capture graded and subjective assessments of long-form generations quality.
+This subjectivity arises from the multifaceted nature of evaluating long-form outputs, where factors like factuality, coherence, clarity, and comprehensiveness each play a role, potentially introducing variability in judgment Bakker et al.
+(2022).
+We can then measure both the classical notion of calibration error, averaged across different correctness levels, as well as new notions of alignment between the correctness and confidence distributions and their utility in selective prediction (§ 3.4).
+Figure 2 shows an overview of our framework, which consists of three modular components: estimating the target correctness distributions, eliciting confidence distributions from LLMs, and measuring calibration between these distributions.
+Our unified framework offers three key advantages.
+1.
+Generalizability: Our framework applies to both long-form and short-form generation tasks by representing correctness and confidence as distributions, regardless of whether the correctness of task is binary, continuous, subjective, or objective.
+2.
+Flexibility: The framework is evaluation-metric agnostic, allowing the integration of any metric or confidence elicitation method, and can adapt as evaluation methods evolve.
+3.
+Interpretability: It provides a nuanced view of uncertainty, enabling decision-makers to assess confidence across multiple correctness levels, fostering greater transparency and trust in the model’s outputs.
+We leverage our framework to measure calibration for several LLMs on multiple datasets across three long-form QA—ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022)—and one summarization task, CNNDM (Nallapati et al., 2016).
+Our results show that our methods excel over baselines by leveraging the model’s nuanced confidence distribution, stronger LLMs like GPT-3.5 don’t necessarily guarantee better calibration, that various calibration metrics complement each other, and LLMs exhibit better calibration on factoid datasets than more open-ended datasets.
+Furthermore, our analysis highlights that fine-tuning and temperature scaling could enhance calibration.
+Finally, we illustrate a practical application of long-form calibration: employing a cascading strategy (Chen et al., 2023a) for selective answering to optimize the cost-effectiveness of long-form text generation.
+In this approach, an open-source model initially handles queries and, based on its confidence levels—assessed using our system—a more advanced API model is engaged as needed.
+This method ensures cost efficiency while maintaining high-performance levels.
+In summary, our contributions are:
+  • A universal calibration framework for text generation tasks, enhancing LLM evaluation for critical applications.
   • Innovative methods for confidence elicitation and calibration measurement, applied to a variety of LLMs.
-  3. 3.
-    Evidence that calibration can be augmented through model fine-tuning, document incorporation, temperature scaling, and mixed elicitation methods.
-  4. 4.
-    A cost-effective model usage strategy, illustrating the practicality of our calibration system in optimizing LLM deployment.
+  • Evidence that calibration can be improved by model fine-tuning and temperature scaling.
+  • A cost-effective model usage strategy, illustrating the practicality of long-form calibration in optimizing LLM deployment.
 2 Related Work
-Measuring Calibration Calibration (Guo et al., 2017, Minderer et al., 2021), the alignment of a model’s confidence with the actual accuracy of its predictions, has been widely studied in language models, whose probabilities derived from logits are generally found to not be calibrated (Jiang et al., 2020, Kadavath et al., 2022, Chen et al., 2023d).
+Measuring Calibration Calibration (Guo et al., 2017, Minderer et al., 2021)has been widely studied in language models, whose probabilities derived from logits are generally found to not be calibrated (Jiang et al., 2020, Kadavath et al., 2022, Chen et al., 2023d).
 Standard metrics to measure the calibration include Expected Calibration Error (ECE) for confidence-accuracy disparity (Naeini et al., 2015), Brier Score for mean squared prediction-outcome differences, and AUROC for assessing confidence-based correct answer identification (Boyd et al., 2013, Kuhn et al., 2023).
 Selective Accuracy@Coverage measures accuracy within the model’s most confident predictions (Liang et al., 2023, Cole et al., 2023).
-However, these metrics are predominantly based on a binary conception of correctness, rendering them less effective for long-form text generation tasks where correctness is better represented as a distribution over a range of scores.
-Improving Calibration Several studies develop improved methods for estimating confidence levels in model responses to enhance the calibration of LMs, such as training external calibrators (Mielke et al., 2022), fine-tuning models to express confidence (Lin et al., 2022), and using sensitivity to in-context learning as a proxy for confidence (Chen et al., 2023c).
-Given the rise of proprietary LLMs with limited access, confidence elicitation methods that don’t require access to logits (Hu et al., 2023) or embeddings have been extensively studied.
-These methods include verbalization-based, which directly asks the model to output its confidence (Lin et al., 2022), consistency-based, which samples multiple responses and uses consistency as a surrogate for confidence (Kadavath et al., 2022, Kuhn et al., 2023, Cole et al., 2023, Tian et al., 2023a), and the hybrid of both (Xiong et al., 2023, Tian et al., 2023b, Chen and Mueller, 2023).
-However, most existing methods still operate under the binary assumption of answer correctness.
-Consequently, confidence is a score that represents the model’s belief that the answer is all correct, inadequate for the complexity of long-form generation tasks.
-3 Calibration Formulation
-In the following sections, we first formalize the long-form generation calibration problem (§ 3.1), and then we introduce three core components in our calibration framework (Figure 2): the correctness distributions of the answers (§ 3.2), the confidence distributions of the LLM on its answers (§ 4), and the calibration metric to measure the difference between correctness distributions and confidence distributions (§ 5).
+However, these metrics, rooted in a binary notion of correctness, fall short for long-form tasks where correctness spans a range, suggesting a distribution-based approach is more apt.
+Improving Calibration Traditional calibration methods focus on post-processing logits (Guo et al., 2017), but with LLMs generating unbounded text, logits could fall short.
+Thus, extracting better confidence scores (i.e., confidence elicitation) has become crucial for improving calibration.
+These include: verbalization, which directly asks the model to output its confidence (Lin et al., 2022), consistency, which uses the uniformity of multiple responses to gauge confidence (Kadavath et al., 2022, Kuhn et al., 2023, Cole et al., 2023, Chen et al., 2023c, Tian et al., 2023a, Lin et al., 2023), and the hybrid of both (Xiong et al., 2023, Tian et al., 2023b, Chen and Mueller, 2023).
+However, these methods often presume binary answer correctness, offering a singular confidence score that fails to capture the nuanced correctness required for long-form tasks.
+Recent work (Zhang et al., 2024) on long-form generation addresses continuous correctness scores but focuses on aligning an uncertainty score with correctness, rather than on improving calibration.
+3 Long-form Generations Calibration
+This section formalizes the long-form generation calibration problem (§ 3.1), and introduces three core components in our calibration framework (Figure 2): the correctness distributions of the answers (§ 3.2), the confidence distributions of the LLM on its answers (§ 3.3), and the calibration metrics to measure how well these two align (§ 3.4).
 3.1 Formulation
@@
 To measure how calibrated the model is, we need three steps.
-First, we apply an evaluator to get the ground-truth correctness distribution $P_{G_{i}}$ where $G_{i}$ is the random variable that denotes the correctness score in answer $A_{i}$.
-Formally,
-$P_{G_{i}}(x)=\text{Pr}(A_{i}\text{ is }x\%\text{ correct})$ (1)
-for $x\in[0,1]$.
-Second, we leverage a confidence elicitation method to extract confidence distribution $P_{C_{i}}$ from an LLM $\mathcal{M}$, where $C_{i}$ is the random variable that denotes the confidence score of the model $\mathcal{M}$ in answer $A_{i}$.
-We further assume that the confidence scores are normalized across $x$ such that they form a valid distribution, with the same domain as the correctness distribution.
-Formally, for $x\in[0,1]$
-$P_{C_{i}}(x)$ $=\text{Pr}(\mathcal{M}\text{ thinks }A_{i}\text{ is }x\%\text{ correct})$ (2)
-Finally, we use $\mathbb{P_{G}}(\mathcal{D})=\{P_{G_{1}},P_{G_{2}},...,P_{G_{n}}\}$ and $\mathbb{P_{C}}(\mathcal{D})=\{P_{C_{1}},P_{C_{2}},...,P_{C_{n}}\}$ to denote the set of correctness and confidence distributions across the whole dataset, where $n$ is the number of data points.
-In our framework, a model is calibrated if its confidence distribution for its generated response closely matches the ground-truth correctness distribution, as measured by a metric $M(\mathbb{P_{G}}(\mathcal{D}),\mathbb{P_{C}}(\mathcal{D}))$.
-3.2 Groundtruth Correctness
-A key challenge in the above formulation lies in obtaining a reliable measure of ground-truth correctness $G(A_{i})$.
-For short-form QA, correctness can be estimated via simple string comparison metrics Rajpurkar et al.
-(2016), but for long-form answers, the most reliable evaluation is usually via human annotation.
-For automatic evaluation, traditional metrics like BLEU and ROUGE are known to have limitations for evaluating long-form QA Liu et al.
-(2023a), and in our preliminary explorations, we found task-specific metrics like the disambiguation-F1 metric for ASQA Stelmakh et al.
-(2022) to correlate poorly with the quality of responses generated by closed-book LLMs prompted without reference texts.
-Hence, in this paper, we adopt the GPT-4 model to evaluate answers, which has been shown to correlate with human judgments in long-form generation (Liu et al., 2023a).
-We ask GPT-4 to range an answer from 0 (worst) to 5 (best).
-The evaluation template includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
-See Appendix E.1 for details.
-To address potential scoring variability caused by the criteria ambiguity and the inherent uncertainty of LLM, we ask the evaluator to repeatedly evaluate each answer, creating a distribution of scores that more accurately represent the answer’s correctness.
-Specifically, given an LLM evaluator $\mathcal{E}$, $N$ evaluations from it, and a score $x\in[0,1]$:
-$P_{G_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{E}(A_{i})_{j}=x)$ (3)
-When N set to 1, the correctness distribution becomes a point-mass distribution, typically observed in tasks with objective ground-truth correctness.
-4 Confidence Elicitation
-We develop two methods to access the model’s confidence distribution about its own answer.
-The first is to explicitly ask the model to self-evaluate its own answer.
-The second is to implicitly assess the model’s confidence in the answer by sampling multiple responses from it and estimating the consistency between them.
-These methods circumvent the need for logit-based approaches (Jiang et al., 2020), which often have limitations.
-Specifically, the normalization of logits-derived probabilities often leads to a skewed distribution, as most small values cluster at the lower end, reducing the model’s discriminative ability and complicating interpretation.
-Additionally, logit-based methods generally imply an overestimation of confidence (Mielke et al., 2022).
-Finally, the inaccessibility of model logits in many commercial, closed-source platforms restricts the practicality of using logit-based methods in these contexts.
-4.1 Self Evaluation
-We prompt the model to perform self-evaluation in order to obtain its estimated confidence distribution.
-The self-evaluation template closely mirrors the template for correctness distribution but omits the reference answer (see Appendix E.3 for details).
-To get the confidence distribution, we ask the model to repeatedly evaluate the same answer with sampling decoding, with the subsequent scores interpreted as the confidence distribution.
-Formally, given an LLM $\mathcal{M}$, $N$ self-evaluations from it, and a score $x\in[0,1]$:
-$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{M}(A_{i})_{j}=x)$ (4)
-Such a sampling method provides a more authentic reflection of the model’s internal distribution than logits.
-By asking the model to assess an answer multiple times, we capture a range of scores that more accurately represent the model’s confidence, which enhances the reliability of the confidence estimation.
-4.2 Self Consistency
-Another key indicator of model confidence is the consistency among multiple responses a model provides for a given question.
-Formally, given a primary answer $A_{i}$, other $N$ answers $A_{i}^{1}...A_{i}^{N}$ sampled from LLM $\mathcal{M}$, a metric for measuring the similarity between two answers $Sim(\cdot,\cdot)$, and a score $x\in[0,1]$:
-$P_{C_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(Sim(A_{i},A_{i}^{j})=x)$ (5)
-Prior research (Kadavath et al., 2022) indicates that such confidence scores correlate to accuracy when the language model is sufficiently large.
-While it’s straightforward to determine if two short answers are consistent, evaluating them in long-form generation is more complex.
-To adapt self-consistency to long-form generation, we introduce four different variations to measure answer similarity in long-form content.
-1.
-Naive: An additional LLM assesses the similarity between two responses.
-2.
-Split: The first answer is divided into individual sentences, and another LLM evaluates the presence of similar statements in the second answer.
-The similarity score is the average of these sentence comparisons.
-3.
-Claim: Building on the Split method, this variation uses a factual detector to identify factual claims in the first answer and checks for their presence in the second response.
-4.
-Named Entity Recoginition (NER): This method further refines the analysis by comparing the overlap of named entities between two responses.
-For a detailed description of these methodologies and the models used, refer to the Appendix B.
-Each of these methods caters to different scenarios and levels of analysis depth, from the broad-strokes assessment of the naive approach to the detail-oriented NER method.
-The choice of method depends on the specific requirements of the task and the desired balance between comprehensiveness and specificity in the consistency analysis.
-5 Calibration Metric
-We propose three distinct calibration metrics to capture how well-calibrated the model is from different perspectives.
-The first metric Correlation examines the relationship between the expected confidence and correctness scores across the entire dataset, serving as an indicator of the model’s proficiency in ranking answers.
-The second, Wasserstein Similarity, measures the congruence between the confidence and correctness distributions for each instance in the dataset, offering a nuanced view of the model’s ability to mirror the shape and spread of the actual correctness in its confidence distribution.
-The third metric Selective F1 assesses the model’s ability to judiciously choose when to respond, reflecting the utility of the confidence scores in picking out the best responses across the dataset.
-These metrics, detailed in the following sections, provide diverse insights into calibration, enriching our understanding of LLM reliability from multiple angles.
-5.1 Correlation
-We assess the dataset-level calibration by measuring the correlation between expected values of confidence and correctness distributions.
-For each answer $A_{i}$ within the dataset, we can calculate the expected confidence score $E(C_{i})$ and the expected correctness score $E(G_{i})$.
-In the whole dataset, we can get a list of expected confidence scores $\mathbf{E_{C}}$ and a list of expected correctness scores $\mathbf{E_{G}}$.
+First, we apply an evaluator to get the target correctness distribution $P_{T_{i}}$ where $T_{i}$ is the random variable that denotes the correctness score in answer $A_{i}$.
+$P_{T_{i}}(x)=\text{Pr}(A_{i}\text{ is }s\text{ correct})$ (1)
+for $s\in\mathcal{S}$, where $\mathcal{S}$ is the space of correctness levels (e.g., normalized from ordinal scores ranging from 0 to 5).
+It should be noted that $s$ could theoretically be a continuous value in the range $[0\%,100\%]$.
+However, since humans tend to make more accurate judgments using discrete ratings due to “rounding bias” (Honda et al., 2022), we approximate continuous correctness with ordinal scores in practice.
+Second, We use a confidence elicitation method to derive the confidence distribution $P_{C_{i}}$ from LLM $\mathcal{M}$, where $C_{i}$ represents the model’s confidence in its answer $A_{i}$.
+We ensure these confidence scores are normalized to form a valid distribution, matching the domain of the correctness distribution.
+For $s\in\mathcal{S}$,
+$P_{C_{i}}(x)=\text{$\mathcal{M}$'s confidence that {A} is {s} correct}$ (2)
+Finally, we design metrics to measure the alignment between $P_{C_{i}}$ and $P_{T_{i}}$ across the dataset.
+3.2 Correctness Distribution Estimation
+To establish correctness distributions as alignment targets, we need to adopt long-form evaluation metrics that integrate aspects such as relevance, coherence, factuality, and helpfulness.
+Traditional metrics like BLEU and ROUGE fail to capture semantic meaning Liu et al.
+(2023a), while factuality-based metrics like FactScore Min et al.
+(2023) may neglect question relevance.
+GPT-4 metrics have gained popularity Li et al.
+(2024) due to their adaptability and comparative accuracy (Jain et al., 2023, Liu et al., 2023a).
+These metrics allow for the integration of various user-prioritized aspects by adjusting evaluation rubrics, providing a balanced approach to both referenced and divergent answers.
+However, they also have limitations like a bias toward longer outputs Zheng et al.
+(2023).
+Our framework is evaluator-agnostic, allowing us to use any correctness distribution metric.
+To identify a practical and actionable metric that best guides human decision-making, we conduct human evaluations to determine alignment with human judgment.
+According to the results in Appendix C, we determined that the GPT-4 metric, with a higher correlation to human judgments compared to the task-specific metric, is more effective for three datasets (e.g., 76.2 v.s 47.8 in ASQA), while the task-specific metric is preferred for another dataset.
+Our framework’s modular design allows replacements of evaluation metrics based on user needs or new developments.
+3.3 Confidence Distribution Elicitation
+There are two common strategies to develop confidence in model responses: explicitly asking the model to verbalize its confidence or implicitly estimating it through self-consistency.
+However, the single confidence score provided by prior studies is ambiguous and non-interpretable in long-form calibration.
+Therefore, we develop two methods tailored for long-form calibration accordingly.
+Continuous Self Evaluation (CSE) We prompt the model to repeatedly perform self-evaluations, where the resulting scores (typically ordinal scores from 0 to 5) are normalized to a [0,1] range and interpreted as a confidence distribution.
+The self-evaluation template closely mirrors the template for correctness evaluation but omits the reference answer (see Appendix F.3 for details).
+Formally, given an LLM $\mathcal{M}$, $N$ self-evaluations:
+$P_{C_{i}}(s)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{M}(A_{i})_{j}=s)$ (3)
+for score $s\in\mathcal{S}$, where $\mathcal{S}$ is the space of correctness levels.
+Such a sampling method provides a more authentic reflection of the model’s internal distribution than logits (Cole et al., 2023).
+By asking the model to assess an answer multiple times, we capture a range of scores that better represent the model’s confidence, which enhances the reliability of the confidence estimation.
+Pairwise Self Consistency (PSC) Another key indicator of model confidence is the consistency among multiple responses a model provides for a given question.
+Given a primary answer $A_{i}$, other $N$ answers $A_{i}^{1}...A_{i}^{N}$ sampled from LLM $\mathcal{M}$, a metric for measuring the similarity between two answers $Sim(\cdot,\cdot)$, and a score $s\in\mathcal{S}$:
+$P_{C_{i}}(s)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(Sim(A_{i},A_{i}^{j})=s)$ (4)
+Assessing similarity in long-form answers is more complex than with short responses (Kadavath et al., 2022).
+To address this, we propose four methods for evaluating similarity in long-form content: 1.
+Naive, assessing overall response similarity with an LLM; 2.
+Split, analyzing sentence-level similarity; 3.
+Claim, evaluating claim matching; 4.
+Named Entity Recognition (NER), focusing on named entity overlap.
+These approaches range from broad to detailed analysis, chosen based on task requirements and the desired analysis depth.
+See Appendix B for more details.
+3.4 Calibration Metrics
+We introduce three key metrics to assess model calibration from various angles.
+Expected correctness error with multi-class (ECE-M) measures the alignment between the model’s stated confidence in reaching a particular level of correctness and the actual likelihood that the model performs at that specified level, across the spectrum from 0 to 1.
+Correlation evaluates the alignment between expected confidence and correctness across the dataset, indicating the model’s proficiency in ranking answers.
+Selective F1 measures the utility of confidence scores in identifying the good answers and abstaining from the rest.
+ECE-M The classical notion of calibration relies on an (answer-correctness) pair of random variables $(A,Y)$ $\in\mathcal{A}\times\{0,1\}$, where $\mathcal{A}$ is the answer space.
+An LLM $\mathcal{M}$ with its confidence elicitation method $h_{\mathcal{M}}$: $\mathcal{A}\rightarrow[0,1]$ is said to be well-calibrated if $\text{Pr}(Y=1|h_{\mathcal{M}}(A)=q)=q$ for $q\in[0,1]$.
+To measure if this holds, traditional ECE(h) Gupta and Ramdas (2021) is: $\mathbb{E}_{A}\left[\left|\text{Pr}(Y=1|h_{\mathcal{M}}(A))-h_{\mathcal{M}}(A)\right|\right]$
+In long-form calibration where the answer correctness is a continuum $Y\in[0,1]$, the probabilistic confidence predictor $h_{\mathcal{M}}$ should predict confidence about each level of correctness and therefore denoted as $h_{\mathcal{M}}$: $\mathcal{A}\times[0,1]\rightarrow[0,1]$.
+The notion of long-form calibration is $\text{Pr}(Y=s|h_{\mathcal{M}}(A,s)=q_{s})=q_{s}$ for every $s\in[0,1]$ and $q_{s}\in[0,1]$ We define ECE-M as the aggregation of ECE scores for all correctness levels.
+In practice, we use discrete levels $s\in\mathcal{S}$ (e.g., ratings from 0-5) for the correctness scores.
+Hence, we calculate an ECE$(s,h)$ conditioned on each $s$,
+$\mathbb{E}_{X}\left[\left|\text{Pr}(Y=s|h_{\mathcal{M}}(A,s))-h_{\mathcal{M}}(A,s)\right|\right]$ (5)
+Then the final ECE-M score is weighted by the frequency of each class:
+[…]
+Correlation While ECE-M focuses on measuring calibration at each correctness level independently, it doesn’t account for the distance between these levels.
+To address this, we also measure the correlation between the expected values of the confidence and correctness distributions for a more comprehensive assessment.
+For each answer $A_{i}$ within the dataset, we can calculate the expected correctness score of the model $E|C_{i}|=\sum_{s}P_{C_{i}}(s)\times s$ and the expected correctness score of the target $E|T_{i}|=\sum_{s}P_{T_{i}}(s)\times s$.
+In the whole dataset, we can get a list of expected correctness scores $\mathbf{E_{C}}$ from model confidence and a list of target expected correctness scores $\mathbf{E_{T}}$.
 Then we can measure the correlation between them.
-$\rho(\mathcal{D})=Corr(\mathbf{E_{C}},\mathbf{E_{G}})$ (6)
-where $Corr(\cdot)$ represents the correlation function (Pearson correlation in our experiments).
-This correlation provides a clear indicator of how well our model’s confidence aligns with its actual correctness across the entire dataset.
-5.2 Wasserstein Similarity
-Wasserstein distance represents how much “work” it takes to transform one distribution into another.
-Wasserstein similarity $S_{W}$ adapts the concept of Wasserstein distance $D_{W}$ in a way that it represents similarity rather than distance between two probability distributions.
-Formally, the instance-level Wasserstein similarity for an answer $A_{i}$ is:
-$S_{W}(A_{i})=1-\frac{D_{W}(G_{i},C_{i})}{d_{max}}$ (7)
-where $d_{max}$ is the maximum distance in the metric space. And the dataset-level calibration score can be obtained by averaging the instance-level calibration score.
-$S_{W}(\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{i}{S_{W}(A_{i})}$ (8)
-5.3 Selective F1
-In selective answering (e.g., Kamath et al., 2020), models only respond when confident about their accuracy, enhancing their reliability.
-To assess the effectiveness of LMs in selective answering, we propose the selective F1 metric ($F1_{\tau_{s}}$).
-This metric quantifies the model’s aptitude in filtering out answers that meet or exceed a predefined correctness threshold, denoted as $\tau_{s}$.
-Alongside this, we define two critical components: selective precision ($P_{\tau_{s}}$) and selective recall ($R_{\tau_{s}}$).
-Selective precision calculates the proportion of selected answers that surpass the correctness threshold $\tau_{s}$ relative to the total number of selected answers.
-In contrast, selective recall compares the number of selected answers meeting this criterion against the total number of correct answers in the dataset that exceed the threshold $\tau_{s}$.
+$\rho(\mathcal{D})=Corr(\mathbf{E_{C}},\mathbf{E_{T}})$ (7)
+where $Corr(\cdot)$ represents the correlation function. This correlation provides a clear indicator of how well our model’s confidence aligns with its actual correctness across the entire dataset.
+Selective F1 In selective answering (e.g., Kamath et al., 2020), models only respond when confident about their accuracy to improve reliability.
+Traditional metrics for selective answering include accuracy@coverage and coverage@accuracy (Tian et al., 2023b), which measures a model’s precision and recall in selecting completely correct answers.
+Similarly, in long-form calibration, it is crucial to assess the model’s selection of answers that are at least "scorrect" using both precision and recall.
+Therefore, we propose the selective F1 metric ($F1_{\tau_{s}}$) to quantify the model’s aptitude in filtering out answers that meet or exceed a predefined correctness threshold $\tau_{s}$.
 Our approach utilizes a dual-threshold system, consisting of a confidence threshold ($\tau_{c}$) and the correctness threshold ($\tau_{s}$), allowing the model to answer questions only if its confidence in the answer’s expected correctness score exceeding $\tau_{s}$ surpasses $\tau_{c}$.
-To more formally define the selective F1, let $\mathcal{A}=\{A_{1},A_{2},...,A_{n}\}$ denote the total set of model’s answers in the dataset and $\mathcal{A}^{*}=\{A_{i}\in\mathcal{A}\mid\text{if }\sum_{s\geq\tau_{s}}C_{i}(s)\geq\tau_{c}\}$ denote the set of selected answers.
-Let the indicator function $I_{\tau_{s}}(A_{i})$ indicate if the expected correctness score $E[G_{i}]$ of $A_{i}$ exceeds $\tau_{s}$
-$I_{\tau_{s}}(A_{i})=\begin{cases}1&\text{if }E[G_{i}]\geq\tau_{s}\\ 0&\text{Otherwise}\end{cases}$ (9)
-The selective precision on the dataset $\mathcal{D}$ is:
+Formally, let $\mathcal{A}=\{A_{1},A_{2},...,A_{n}\}$ denote the total set of model’s answers in the dataset and $\mathcal{A}^{*}=\{A_{i}\in\mathcal{A}\mid\sum_{s\geq\tau_{s}}P_{C_{i}}(s)\geq\tau_{c}\}$ denote the set of selected answers.
+Let the indicator function $I_{\tau_{s}}(A_{i})$ indicate if the expected correctness score $E[T_{i}]$ of $A_{i}$ exceeds $\tau_{s}$:
+$I_{\tau_{s}}(A_{i})=\begin{cases}1&\text{if }E[T_{i}]\geq\tau_{s}\\ 0&\text{Otherwise}\end{cases}$ (8)
+The selective precision $P_{\tau_{s}}$ on the dataset $\mathcal{D}$ is the proportion of selected answers that surpass the correctness threshold $\tau_{s}$ relative to the total number of selected answers:
 $P_{\tau_{s}}(\mathcal{D})=\frac{\sum_{A_{i}\in\mathcal{A}^{*}}I_{\tau_{s}}(A_{i})}{|\mathcal{A}^{*}|}$ (9)
-The selective recall is:
+The selective recall $R_{\tau_{s}}$ compares the number of selected answers meeting this criterion against the total number of correct answers in the dataset that exceed the threshold $\tau_{s}$:
 $R_{\tau_{s}}(\mathcal{D})=\frac{\sum_{A_{i}\in\mathcal{A}^{*}}I_{\tau_{s}}(A_{i})}{\sum_{A_{i}\in\mathcal{A}}I_{\tau_{s}}(A_{i})}$ (10)
-The selective F1 takes both recall and precision into account:
+The selective F1 combines recall and precision:
 $F1_{\tau_{s}}(\mathcal{D})=2\frac{P_{\tau_{s}}(\mathcal{D})R_{\tau_{s}}(\mathcal{D})}{P_{\tau_{s}}(\mathcal{D})+R_{\tau_{s}}(\mathcal{D})}$ (11)
-ASQA ELI5
-Model Method Corr Wass $F1_{0.8}$ Score Corr Wass $F1_{0.8}$ Score
-Llama2-13b Self Eval 18.1 71.5 51.0 51.3 21.8 75.0 39.6 53.8
-Self Cons 48.0 75.8 3.6 28.8 61.0 39.8
-Llama2-70b Self Eval 16.4 64.9 60.3 59.4 18.8 78.9 54.9 61.7
-Self Cons 44.3 77.0 29.7 15.5 71.4 51.4
-Vicuna-13b Self Eval 30.2 70.8 30.0 50.8 31.2 76.6 36.7 53.0
-Self Cons 58.2 72.0 37.6 35.3 63.8 36.9
-GPT-3.5-turbo Self Eval 3.6 74.3 83.7 72.6 26.2 76.9 63.5 63.0
-Self Cons 30.5 77.7 58.4 18.5 68.6 55.2
-QAMPARI CNNDM
-Model Method Corr Wass $F1_{0.4}$ Score Corr Wass $F1_{0.8}$ Score
-Llama2-13b Self Eval 25.7 66.6 29.0 13.3 19.5 96.5 92.0 77.0
-Self Cons 46.0 71.6 42.3 9.6 73.2 24.5
-Llama2-70b Self Eval 18.6 41.9 30.1 14.6 6.4 96.9 93.6 77.6
-Self Cons 17.1 61.0 39.1 22.7 72.4 20.2
-Vicuna-13b Self Eval 16.2 62.6 27.0 11.4 3.9 80.9 64.2 78.2
-Self Cons 49.5 78.7 39.7 18.9 78.5 41.7
-GPT-3.5-turbo Self Eval 28.7 57.6 53.1 24.0 17.0 96.2 94.7 78.2
-Self Cons 44.4 65.8 56.0 18.9 77.8 32.1
-Table 1: Calibration Performance Across Four Tasks (in %): “Corr” for Correlation, “Wass” for Wasserstein Similarity, “$F1_{\tau_{s}}$” for Selective F1 Score at threshold $\tau_{s}$, and “Score” representing expected correctness.
-“Self Eval” for self-evaluation and “Self Cons” for self-consistency.
-Bold indicates the best score.
-Key insights include: 1) Higher model capability does not ensure better calibration; 2) Models may exhibit calibration on certain metrics but not others; 3) On factoid datasets (ASQA, QAMPARI), self-consistency outperforms self-evaluation.
+In our experiments, we select $\tau_{s}$ as the nearest correctness level greater than the best LM’s average correctness score.
+For $\tau_{c}$, we choose the value that yields the highest selective F1 score on the development split.
+ASQA QAMPARI ELI5 CNNDM
+Method ECE-M Corr $F1_{0.8}$ ECE-M Corr $F1_{0.4}$ ECE-M Corr $F1_{0.8}$ ECE-M Corr $F1_{0.8}$
+SL 28.2$\uparrow$ 0.7$\uparrow$ 0.0 $\uparrow$ 27.4$\uparrow$ 10.9$\uparrow$ 8.0$\uparrow$ 29.6$\uparrow$ -11.9$\uparrow$ 0.0$\uparrow$ 77.2$\uparrow$ -7.5$\uparrow$ 0.0$\uparrow$
+BSE 32.8$\uparrow$ 14.2$\uparrow$ 57.6 $\uparrow$ 25.2$\uparrow$ 16.8$\uparrow$ 33.2$\uparrow$ 30.3$\uparrow$ 11.7$\uparrow$ 46.7$\uparrow$ 78.5$\uparrow$ 11.2$\uparrow$ 90.6$\uparrow$
+CSE 29.0$\downarrow$ 16.3$\uparrow$ 58.5 $\uparrow$ 42.8$\uparrow$ 21.9$\uparrow$ 33.4$\uparrow$ 31.2$\uparrow$ 26.9$\uparrow$ 48.2$\uparrow$ 15.2$\downarrow$ 19.2$\uparrow$ 92.0$\uparrow$
+ASC 35.9$\uparrow$ 27.1$\uparrow$ 5.2 $\uparrow$ 46.0$\uparrow$ 38.6$\uparrow$ 38.5$\uparrow$ 38.4$\uparrow$ 16.7$\uparrow$ 7.9$\uparrow$ 63.2$\uparrow$ 8.8 $\uparrow$ 44.2$\uparrow$
+PSCF1 28.8$\uparrow$ 27.1$\uparrow$ 33.5 $\uparrow$ 38.4$\uparrow$ 38.6$\uparrow$ 42.7$\uparrow$ 27.1$\uparrow$ 16.7$\uparrow$ 20.7$\uparrow$ 57.1$\uparrow$ 8.8$\uparrow$ 79.5$\uparrow$
+PSC 18.3$\downarrow$ 46.8$\uparrow$ 61.6 $\uparrow$ 26.2$\downarrow$ 39.1$\uparrow$ 44.0$\uparrow$ 24.9$\downarrow$ 24.9$\uparrow$ 46.2$\uparrow$ 64.5$\uparrow$ 15.5$\uparrow$ 90.0$\uparrow$
+Table 1: Calibration Performance Comparison Among Different Confidence Elicitation Methods Across Four Tasks (in %): “ECE-M” for expected correctness error with multi-class, “Corr” for Correlation, “$F1_{\tau_{s}}$” for Selective F1 Score at threshold $\tau_{s}$.
+Results represent averages from five models.
+Methods with * served as baselines.
+For “Corr”, “$F1_{\tau_{s}}$”, and “Score”, $\uparrow$ means better than corresponding baseline while $\downarrow$ is worse.
+For “ECE-M”, $\downarrow$ is better while $\uparrow$ is worse.
+The best score among all confidence elicitation methods is bolded.
+Key insights: 1) Self-Consistency (PSC) outperforms Self-Evaluation on factoid datasets; 2) Our methods PSC and CSE surpass baselines; 3) Different metrics offer complementary insights
 4 Experiments and Results
 4.1 Setup
-Models and Data We measure different sized LLMs’ calibration, including Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), vicuna-13b (Zheng et al., 2023), GPT-3.5-turbo, acorss three long-form QA tasks: ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022), and one summarization task: CNNDM (Nallapati et al., 2016).
-Details of each dataset can be found in Appendix A.
-Correctness Evaluation We apply GPT-4 to evaluate the model-generated answers to approximate the ground-truth correctness distributions for ASQA, ELI5, and CNNDM.
-In QAMPARI, given that the answer is a list of entities, we evaluate using the F1-5 metric (Gao et al., 2023), calculating the F1 score by the exact match with the gold answer list and defining 100% recall for predictions with at least 5 correct answers.
-In preliminary experiments, we found that samples from the GPT-4 evaluator had very low variance in terms of the final score assigned to the answers, hence in the following results we only use a single sample, resulting in a point mass distribution for correctness.
-Confidence Elicitation We contrast self-evaluation and self-consistency methods for confidence assessment across various tasks, employing task-specific strategies to compare similarities.
-For ASQA, where answers hinge on factual claims, we utilize a self-consistency approach focused on claims.
-In the case of QAMPARI, which involves listing entities, we adopt self-consistency-NER.
-For the ELI5 dataset, a more general self-consistency-naive strategy is applied due to the broad nature of the questions.
-Lastly, for CNNDM, where summaries are composed of key sentences, we implement a self-consistency-split for detailed analysis.
-These tailored approaches ensure the most effective evaluation of each dataset’s unique characteristics.
+Models and Data We measure different sized LLMs’ calibration, including Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Vicuna-13b (Zheng et al., 2023), Llama-3-8b-Instruct, GPT-3.5-turbo, across three long-form QA tasks: ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022), and one summarization task: CNNDM (Nallapati et al., 2016).
+Details of datasets can be found in Appendix A.
+Correctness Evaluation We apply GPT-4 to evaluate target correctness distributions for ASQA, ELI5, and CNNDM.
+In QAMPARI where the answer is a list of entities, we evaluate using the F1-5 metric, calculating the F1 score by the exact match with the gold answer and defining 100% recall for predictions with at least 5 correct answers.
+4.2 Confidence Elicitation Methods
+In addition to our methods CSE and PSC (see Appendix D.3 for similarity measurement choosing), we established baselines for self-evaluation, self-consistency, and logits-based approaches.
+This is because prior studies lack directly applicable baselines, primarily due to the non-interpretable nature of single confidence scores.
+Sentence Likelihood (SL): Based on prior studies using logits to gauge model confidence, we adopt sentence likelihood as a baseline measure, which typically results in a confidence distribution focused at the lowest score in long-form answers.
+Binary Self-Evaluation (BSE): following previous work Kadavath et al.
+(2022) that asks model to self-evaluate if its answer is true several times, using the frequency of true as model’s confidence score towards the answer being true.
+Then we adapt such a single score as a distribution focus solely on the values 0 and 1.
+Average Self-Consistency (ASC): Following prior work Xiong et al.
+(2023) using the average consistency between these candidate responses and the original answer then serves as a single measure of confidence score, we adopt the simple f1-token score to measure the consistency to adapt it to long-form generations.
+Then we treat the single score as a point mass distribution.
+Pairwise Self-Consistency F1 (PSC-F1): Still using F1 to measure the consistency like ASC, but we directly treat the pairwise consistency scores as a distribution without aggregating, thereby keeping model intrinsic understanding about different correctness levels.
 4.3 Main Results
+In Table 1, we evaluate the calibration performances of various confidence elicitation methods by averaging the scores across all models.
+Table 1 shows that our methods, CSE and PSC, generally outperform their respective baseline categories and also surpass the logits-based method SL.
+Key findings from the results include:
+Self-Consistency Outperforms Self-Evaluation on Factoid Datasets Self-consistency methods typically outperform self-evaluation on factoid datasets like ASQA and QAMPARI.
+However, their effectiveness diminishes in more subjective tasks such as ELI5 or CNNDM.
+We hypothesize that this is because self-consistency is more readily quantifiable in factoid datasets, where the agreement between answers can be assessed based on factual consistency, thus providing clearer criteria.
+Conversely, in open-ended datasets, the consistency between answers is more ambiguous, making it more difficult to measure.
+Nuanced Self-Evaluation Enhances Calibration CSE generally outperforms BSE by providing detailed confidence estimates at each correctness level.
+However, the overall improvement remains constrained by the intrinsic limitations of LMs in self-evaluating their correctness, which sometimes hampers accurate estimations.
+Pair-wise Similarities Distribution and Task-tailored Similarity Measurement Help Calibration Both ASC and PSC-F1 measure similarity with token-level F1.
+However, PSC-F1 treats these scores as a distribution rather than averaging them, leading to better ECE-M and selective F1.
+PSC further enhances calibration by adopting a task-specific, detailed measurement of similarity, outperforming FSC in all four tasks.
+ASQA QAMPARI
+Model ECE-M Corr $F1_{0.8}$ Score ECE-M Corr $F1_{0.4}$ Score
+Llama2-13b 15.9 48.0 47.5 51.3 30.5 46.0 42.3 13.3
+Llama2-70b 14.7 44.3 61.9 59.4 29.3 17.1 37.0 14.6
+Vicuna-13b 20.1 58.2 56.8 50.8 14.4 49.5 42.6 11.4
+Llama-3-8b 14.3 53.2 65.1 54.9 33.0 38.3 42.0 14.1
+GPT-3.5-turbo 26.7 30.5 76.7 72.6 23.7 44.4 56.2 24.0
+ELI5 CNNDM
+Model ECE-M Corr $F1_{0.8}$ Score ECE-M Corr $F1_{0.8}$ Score
+Llama2-13b 36.0 21.8 40.1 53.8 12.6 19.5 92.0 77.0
+Llama2-70b 32.8 18.8 54.4 61.7 13.8 6.4 93.6 77.6
+Vicuna-13b 19.4 31.3 34.9 53.0 32.7 3.9 64.2 78.2
+Llama-3-8b 34.9 36.7 48.4 57.1 9.3 49.0 86.9 77.8
+GPT-3.5-turbo 32.7 26.2 63.4 63.0 7.8 17.0 94.7 78.2
+Table 2: Comparison of Calibration Performance Across Models for Four Tasks (in %): We identify the optimal confidence elicitation method for each task and compare the performance of various models using this method.
+“Score” means the model’s average correctness score on that task.
+A key observation is that more powerful LMs do not necessarily exhibit better calibration, although they tend to perform better in selective answering.
+Calibration metrics complement each other A simplistic approach like SL, which allocates all the probability mass to the point of score 0, can misleadingly show decent ECE-M (28.2%) in specific cases like ASQA.
+However, its negative correlation (0.7%) and zero F10.8 underscore an ineffective confidence distribution.
+Similarly, BSE in CNNDM may achieve a high F10.8 (90.6%) by overestimating answer correctness, but this does not truly reflect response quality (correlation: 11.2%) or provide well-calibrated probabilities across correctness levels, resulting in a bad ECE-M (ECE-M: 78.5%).
+Hence, a comprehensive evaluation using multiple metrics is essential for a balanced assessment of model calibration.
 Larger models are not necessarily better calibrated.
-Table 1 reveals a surprising trend: Vicuna-13b, though performing poorly on ASQA, ELI5 and QAMPARI, has the highest correlation with self-consistency confidence on the same datasets.
-This might be due to the fact that Vicuna-13b is the only model instruction fine-tuned in a supervised fashion, in contrast to the other models using reinforcement learning human feedback.
-RL finetuning has been shown to lead to miscalibration in LMs by biasing generations towards highly rewarded behaviors (Kadavath et al., 2022).
+In Table 2, we focus on the calibration performance of individual models when paired with the best-performing confidence elicitation method for each task.
+Table 2 shows that despite poor performance on ASQA and QAMPARI, Vicuna-13b has the highest correlation across these datasets.
+It might be because reinforcement learning for other models causes miscalibration by encouraging overfitting to rewarded behaviors (Kadavath et al., 2022).
 Scaling the temperature could enhance the calibration of LLMs fine-tuned using RL (see § 4.4).
-Additionally, Llama-2-13b demonstrates a higher correlation than its larger counterpart, Llama-2-70b, in $3$ out of $4$ datasets with both confidence elicitation methods.
+Additionally, Llama-2-13b demonstrates a higher correlation than its larger counterpart, Llama-2-70b.
 However, GPT-3.5-turbo, the strongest model, consistently scores the highest in selective F1 across all datasets.
-Selective F1 metrics blend performance and calibration and tend to yield higher scores for stronger models.
-A model can be calibrated on some dimensions but not on others.
-Calibration metrics offer complementary perspectives.
-A model may exhibit low correlation but high Wasserstein similarity, or vice versa.
-For instance, GPT-3.5-turbo on the ASQA dataset shows low correlation ($3.6$) but high Wasserstein similarity ($74.3$) with self-evaluation confidence.
-In contrast, Vicuna-13b demonstrates the opposite trend on ASQA with self-consistency confidence (correlation: $58.2$, Wasserstein similarity: $72.0$).
-This discrepancy arises because correlation and Wasserstein similarity capture different calibration aspects: correlation reflects the model’s ability to appropriately rank good and bad answers, while Wasserstein similarity evaluates how closely a model’s confidence mirrors actual correctness.
-Self-Consistency is better than Self-evaluation on factoid datasets.
-In factoid datasets (ASQA and QAMPARI), self-consistency confidence elicitation outperforms self-evaluation for all models.
-On QAMPARI, all models show better calibration with self-consistency across all four metrics, while on ASQA, the trend holds for all but selective F1.
-While self-consistency scores can serve as indicators of answer quality, they are not guaranteed to be the same as ground-truth correctness due to the inability to fully grasp the ground-truth criteria.
-However, the effectiveness of self-consistency diminishes in more open-ended tasks like ELI5 or the CNNDM summarization dataset.
-As tasks become more subjective, the power of self-consistency in model calibration begins to wane.
+This performance can be attributed to the model’s ability to generate a larger volume of high-quality answers, increasing the probability of selecting superior responses even if it is not particularly well-calibrated.
+Consequently, the Selective F1 metric blends performance and calibration, and tends to favor more capable models due to their higher output of quality answers.
 4.4 Improving Calibration
-We delve into four strategies aimed at enhancing the calibration score of LLMs: fine-tuning the model to do self-evaluation, retrieving source documents, scaling the temperature, and hybrid confidence elicitation strategies.
-Fine-tuning Our study examines enhancing model calibration through fine-tuning, which is conducted on the ASQA dataset.
-We generate a self-evaluation training dataset using GPT-4 to assess answers from various models to ASQA’s training set questions.
-This dataset comprises inputs of questions and corresponding model answers, with outputs including a score and an explanation for each answer.
-The Llama2-13b model is then fine-tuned using LoRA (Hu et al., 2021) on this dataset, aiming to improve its scoring accuracy and explanatory capabilities.
-We also consider the impact of self-evaluation finetuning on the model’s ability to generate responses, and further experiment with finetuning the model on a combined objective for self-evaluation as well as generating answers from the ASQA training set.
-For comparison, we also fine-tune the model using generation data only.
-See Appendix D.1 for details of the training methodology.
-As Table 2 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task, as well as the limitation of LORA fine-tuning.
+We delve into different strategies to enhance calibration: fine-tuning, scaling the temperature, adding source documents (Appendix E.2), and hybrid confidence elicitation (Appendix E.3).
+Fine-tuning Our study explores three fine-tuning strategies to improve model calibration on the ASQA dataset: fine-tuning the model for self-evaluation (using questions and model answers to produce scores and explanations), fine-tuning the model for generation (generating answers from questions), and a hybrid of both.
+GPT-4 synthesizes self-evaluation data by assessing different models’ answers to questions from the ASQA training set, while the generation data originates from the ASQA training set itself.
+We apply LoRA (Hu et al., 2021) fine-tuning to the Llama2-13b model.
+See Appendix E.1 for experiment details.
+As Table 3 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task and the limitation of LORA.
 Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’).
 The model becomes more confident in terms of self-consistency after fine-tuning.
-[…]
-Self-Evaluation
-None 18.1 $\uparrow$ 71.5 $\uparrow$ 51.0 $\uparrow$ 51.3 $\uparrow$
-Evaluation 13.6 $\downarrow$ 67.0 $\downarrow$ 62.0 $\uparrow$ 49.2 $\downarrow$
-Generation 20.0 $\uparrow$ 72.4 $\uparrow$ 52.8 $\uparrow$ 52.1 $\uparrow$
-Eval + Gen 23.9 $\uparrow$ 74.2 $\uparrow$ 46.6 $\downarrow$ 50.1 $\downarrow$
-Self-Consistency
-None 48.0 $\uparrow$ 75.8 $\uparrow$ 3.6 $\uparrow$ 51.3 $\uparrow$
-Evaluation 46.9 $\downarrow$ 76.7 $\uparrow$ 10.7 $\uparrow$ 49.2 $\downarrow$
-Generation 58.9 $\uparrow$ 76.4 $\uparrow$ 32.4 $\uparrow$ 52.1 $\uparrow$
-Eval + Gen 54.5 $\uparrow$ 78.4 $\uparrow$ 18.9 $\uparrow$ 50.1 $\downarrow$
+Data Corr ECE-M F10.8 Score
+Self-Evaluation (CSE)
+None 18.1 $\uparrow$ 30.3 $\uparrow$ 50.8 $\uparrow$ 51.3 $\uparrow$
+Evaluation 13.6 $\downarrow$ 32.4 $\uparrow$ 52.3 $\uparrow$ 49.2 $\downarrow$
+Generation 20.0 $\uparrow$ 26.2 $\downarrow$ 53.4 $\uparrow$ 52.1 $\uparrow$
+Eval + Gen 23.9 $\uparrow$ 20.2 $\downarrow$ 46.6 $\downarrow$ 50.1 $\downarrow$
+Self-Consistency (PSC)
+None 48.0 $\uparrow$ 15.9 $\uparrow$ 47.5 $\uparrow$ 51.3 $\uparrow$
+Evaluation 46.9 $\downarrow$ 13.6 $\downarrow$ 56.1 $\uparrow$ 49.2 $\downarrow$
+Generation 58.9 $\uparrow$ 14.5 $\downarrow$ 59.5 $\uparrow$ 52.1 $\uparrow$
+Eval + Gen 54.5 $\uparrow$ 12.2 $\downarrow$ 50.7 $\uparrow$ 50.1 $\downarrow$
 Table 3: Comparison among raw and fine-tuned Llama-2-13b on ASQA.
 “None” for the untrained model, “Evaluation” for the model fine-tuned with the self-evaluation dataset, “Generation” for the model fine-tuned with the ASQA generation data, and “Eval+Gen” for the model fine-tuned with the hybrid dataset combined by self-evaluation dataset and generation data.
-Source Documents We examine how additional contextual information influences LLM calibration.
-Using the Llama-2-13b-chat and GPT-3.5-turbo on ASQA dataset, with a focus on self-consistency as our confidence eliciation method, we test the model’s performance when supplemented with different types of source documents.
-We experiment with two types of documents: random documents related to the question and ‘oracle’ documents directly relevant to the answers, as included in the dataset release Stelmakh et al.
-(2022).
-Our findings, illustrated in Table 3, reveal that oracle documents can enhance model performance and calibration across two out of three metrics for both models, while random documents may be less effective.
-These results underscore the importance of relevant contextual information in model calibration.
-[…]
-[…]
-R 39.0 $\downarrow$ 75.6 $\downarrow$ 30.6 $\uparrow$ 59.6 $\uparrow$
-O 49.5 $\uparrow$ 75.6 $\downarrow$ 20.1 $\uparrow$ 65.8 $\uparrow$
-[…]
-R 36.6 $\uparrow$ 71.8 $\downarrow$ 68.2 $\uparrow$ 64.8 $\downarrow$
-O 27.0 $\downarrow$ 79.5 $\uparrow$ 72.8 $\uparrow$ 75.8 $\uparrow$
-Table 3: How source documents affect the calibration score.
-In the document column, “N” means no documents in the input prompt, “R” means randomly selected documents relevant to the topic, and “O” means the oracle documents relevant to the answer.
-$\uparrow$ denotes that the calibration score goes up when adding documents, while $\downarrow$ means going down.
 Temperature We adjust the generation temperature for Llama2-13b from 0.2 to 1 to examine its impact on calibration.
-The result in Figure 3 reveals a consistent upward trend in all calibration metrics.
+The result in Figure 3 reveals consistent improvements in all calibration metrics.
 Notably, the model’s performance initially improves and then deteriorates.
-This observation implies that modulating the generation temperature can enhance the calibration of the reinforcement-learned model, which is consistent with previous findings observed in short-form answer generation (Kadavath et al., 2022).
+This observation implies that modulating the generation temperature can enhance the calibration of the model.
 […]
@@
 Figure 3: Calibration varies with temperature scaling.
-Hybrid Confident Elicitation Prior work (Xiong et al., 2023, Tian et al., 2023b) found that combining self-evaluation and self-consistency can yield a more accurate confidence score.
-In our study, we explore whether this finding applies to long-form generation.
-Specifically, for an answer $A_{i}$, we merge confidence distributions obtained through both self-evaluation ($C^{eval}_{i}$) and self-consistency ($C^{consist}_{i}$).
-The resultant hybrid confidence distribution, $C^{hybrid}_{i}=\alpha C^{eval}_{i}+(1-\alpha)C^{consist}_{i}$, is modulated by the parameter $\alpha$, allowing us to adjust the influence of each distribution.
-Our experiments are conducted on the ASQA dataset and utilize the correlation as the calibration metric.
-As shown in Figure 4, we observe that the correlation between confidence and correctness initially increases but then declines as $\alpha$ is varied from 0 to 1.
-However, this trend is not observed in other metrics (See Figure 7), indicating that while hybrid calibration elicitation may enhance calibration in terms of correlation, it may not have the same impact on other dimensions.
-[…]
-Figure 4: Correlation vs.
-Weight Variation for hybrid confidence elicitation.
-$\alpha$ is the weight for self-evaluation confidence.
-$\alpha=0$ for solely self-consistency and $\alpha=1$ for solely self-evaluation.
 4.5 Application
-We showcase a cost-effective application of our system (illustrated in Figure 5) by using a cascade of language models with varying capabilities for selectively responding to queries, aiming for optimal performance within a limited API budget.
-We initially employ a set of calibrated open-source LMs (selected by our calibration framework) to handle queries.
-Complex queries, identified by the lower confidence of these models (derived from our self-consistency confidence elicitation method), are escalated to more advanced API LMs.
-This tiered approach strikes a balance between cost-efficiency and performance in LLM applications.
+We showcase an application of long-form calibration in Figure 4: a cost-effective cascading strategy using language models of varying capabilities to efficiently handle queries within an API budget constraint.
+Initially, an open-source model (Llama-2-13b in our experiment) address questions where it believes the answer has a probability higher than $\tau_{c}$ that the answer’s correctness score is above $\tau_{s}$.
+Complex queries, flagged by lower model confidence, are escalated to a more advanced API LM (GPT-4).
+Adjusting $\tau_{c}$ between 0 and 1 controls how many queries reach GPT-4, balancing answer quality with API budget constraints.
+We benchmark using the open-source LM for a zero API budget and the commercial LM for full-budget scenarios.
 Figure 4: The illustration of LLM Cascade.
-Our experiments, conducted on the ASQA and QAMPARI datasets, involve cascading a series of calibrated open-source models, culminating with GPT-3.5-turbo.
-See Appendix D.3 for model selection and cascading details.
-We begin with the first open-source model, addressing questions where it believes the answer has a probability higher than $\tau_{c}$ that the answer’s correctness score is higher than $\tau_{s}$=0.6.
-Subsequent questions are then processed through the remaining open-source models with the same method.
-Any queries not confidently handled by these models are ultimately directed to GPT-3.5-turbo for resolution.
-By adjusting $\tau_{c}$ from 0 to 1, we can vary the number of queries that will be sent to GPT-3.5-turbo, allowing us to analyze the quality of the final answer as the budget increases for calling the API.
-We use the strongest open-source model’s results as a baseline for a zero API budget, and GPT-3.5-turbo’s results for a full budget scenario.
-As a baseline, we compare to a strategy where, within a limited number of API requests, a random selection of queries is addressed by GPT-3.5-turbo and the rest are handled by the strongest open-source model.
-For each API budget, we randomly select questions for GPT-3.5-turbo using 10 different random seeds and calculate the mean and std for the results.
-As Figure 6 shows, our LLM cascade method always outperforms the random response strategy across various budget levels.
-We don’t consider the API cost for obtaining self-consistency confidence scores from open-source models, since GPT-3.5 could be substituted with other open-source NLI models for verifying sentence presence in the second answer.
-000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111606060606565656570707070BudgetCorrectnessASQA
-000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111151515152020202025252525BudgetCorrectnessQAMPARI
-Figure 6: Variation in Expected Correctness Score with API Budget Allocation on ASQA and QAMPARI: This graph contrasts the LLM Cascade approach (blue) with the Random Response strategy (red).
-The red line shows the mean correctness score from 10 random seeds, while the light red shaded area indicates the range of standard deviation.
-Notably, a budget of 1 corresponds to the correctness achieved by GPT-3.5-turbo, while a budget of 0 reflects the performance of the best-performing open-source model selected for LLM cascade.
+Our experiments utilize the ASQA and QAMPARI datasets to evaluate four distinct confidence elicitation strategies: PSC, ASC, CSE, and BSE.
+Additionally, we incorporate a baseline strategy where, under a constrained number of API requests, a random selection of queries is processed by Llama-2-13b, with the remaining handled by GPT-4.
+For each API budget scenario, questions are randomly assigned to Llama-2-13b using 10 different random seeds, and we calculate the mean and standard deviation of the results.
+We focus on the success rate, which we define as the percentage of answers that meet or exceed a user-specified score threshold.
+This metric is reported both for the overall dataset and for the subset of queries selected and handled by Llama-2-13b, illustrating both the general effectiveness of our cascading model and the selective answering capabilities of the individual model.
+As shown in Figure 5, PSC generally outperforms the other methods, with CSE and BSE yielding comparable results that follow.
+ASC, in contrast, performs the poorest, comparable to the random selection strategy.
+These results highlight the pivotal role of advanced calibration techniques, confidence elicitation methods in our case, in boosting the practical utility and cost-efficiency of LLMs when API usage is limited.
+(a) ASQA Overall
+(b) ASQA Selected
+(c) QAMPARI Overall
+(d) QAMPARI Selected
+Figure 5: Variation in Success Rate by API Budget Allocation on the ASQA and QAMPARI Datasets for All Queries and Those Selected by Llama-2.
 5 Conclusion
-In conclusion, our research introduces a comprehensive calibration system designed to evaluate LLMs across both short-form and long-form text generation.
-We have developed confidence elicitation methods that derive confidence distributions from LLMs.
-To assess calibration, we introduced three new metrics, each tailored to capture different aspects of LLM performance.
-Our findings reveal that larger LLMs do not necessarily equate to better calibration, and the degree of calibration significantly varies across different datasets.
-This challenges prevailing assumptions in the field and underscores the need for nuanced approaches in LLM development and assessment.
-Furthermore, our study explores various strategies to enhance LLM calibration, offering practical solutions to common issues faced in the field.
-We also present a potential application of our system, demonstrating its effectiveness in optimizing LLMs performance within the constraints of limited API budgets.
-The system and tools we present are crucial for further improving the liability of LLMs.
-Our research highlights a couple of research directions, including (1) Advancing Calibration Techniques.
-Develop more sophisticated methods like integrating human feedback into the calibration process, incorporating machine learning algorithms that can adaptively adjust calibration based on the context or type of query, and so on.
-(2) Cross-Domain Applications.
-Extend the application of our system to other domains such as medical text generation, legal document analysis, education chatbots and so on.
-(3) Real-time Calibration.
-Develop methods for real-time calibration of LLMs in dynamic environments (e.g.
-interactive chatbots, updated knowledge base).
-(4) Pre-training with Calibration.
-Incorporate calibration into the pre-training phase of LLMs.
-(5) Ethical and Biases Consideration.
-Investigating how different calibration strategies affect the fairness and bias in LLM outputs can be crucial for responsible AI.
-(6) Scalability and Efficiency.
-Explore ways to improve the scalability and efficiency of our calibration system, making it applicable to large-scale, real-world scenarios.
+Our study presents a novel calibration system for evaluating LLMs in long-form generation.
+Our results challenge the assumption that larger LLMs are always calibrated better and show calibration variability across datasets.
+Additionally, we propose methods to improve LLM calibration and show an application that optimizes performance under API budget constraints.
+The system we present is crucial for further improving the liability of LLMs.
+6 Limitation
+Our study faces three primary limitations.
+First, we rely on GPT-4 to estimate the target correctness distribution.
+But as tasks become more subjective, consensus on humans’ evaluations may decrease.
+This wider target correctness distribution becomes challenging for both GPT-4 and human annotators to accurately capture.
+This limitation is inherent to natural language generation (NLG) evaluation and lies beyond the purview of our project.
+Our framework operates under the premise that a target correctness distribution exists and concentrates on calibration which aligns the model’s confidence with this assumed target.
+Second, our experiments focus on long-form QA and do not extend to specialized domains such as law, medicine, or education, where the calibration of LLMs could have significant real-world implications.
+Lastly, our self-consistency method is computationally intensive, posing a challenge for practical applications.
+There is a need for more efficient approaches in real-world settings.
+Replicability:
+Codes: https://github.com/kkkevinkkkkk/calibration
 Appendix A Dataset
@@
 (2023b); Zhang et al.
-(2023a) has found that the provided reference summaries are not of very good quality and zero-shot LLMs summaries are preferred by human annotators over the reference summaries.
+(2023) has found that the provided reference summaries are not of very good quality and zero-shot LLMs summaries are preferred by human annotators over the reference summaries.
 In the ACLE (Automatic LLMs’ Citation Evaluation) (Gao et al., 2023), a pioneering benchmark for assessing LLMs’ citation capabilities, a subset of 1,000 examples is randomly selected from the development sets of ASQA, ELI5, and QAMPARI to form a test set for each task.
@@
 Instead, our approach aims to secure specific consistency ratings that reflect the model’s assurance in its primary answer, offering a general overview of the answers’ similarity.
+The template for similarity measuring can be found in Appendix F.4.
 Sentence Split For a more detailed similarity analysis between the two answers, we split the first answer into individual sentences.
@@
 Some may be filler or less informative, potentially skewing the similarity assessment.
+The template can be found in Appendix F.4.
 Claim To further refine the approach, we focus on sentences that make factual claims.
@@
 We use a Roberta-large (Liu et al., 2019) trained with SpanMarker framework  [^3] https://github.com/tomaarsen/SpanMarkerNER, which can be replaced by other NER models.
-Appendix C Main Experiment
+Appendix C Evaluation Metric
+C.1 GPT-4 metric
+We ask GPT-4 to range an answer from 0 (worst) to 5 (best), which is then normalized to [0, 1].
+See Appendix F.1 for details of the template.
+To mitigate scoring variability from criteria ambiguity and LLM uncertainty, we have evaluators repeatedly score each answer, forming a score distribution that better reflects its correctness.
+Specifically, given an LLM evaluator $\mathcal{E}$, $N$ evaluations from it, and a score $x\in[0,1]$, the correctness distribution is:
+$P_{T_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{E}(A_{i})_{j}=x)$ (12)
+C.2 Human Evaluation
+We utilize GPT-4 to assess answers across different tasks, including ASQA, ELI5, and CNNDM, using GPT-4 scores as a proxy for the target distribution.
+To demonstrate the better alignment of GPT-4 scores with human preferences over task-specific metrics, we focus our human evaluation efforts on the long-form tasks of ASQA and ELI5.
+This approach is supported by prior research indicating GPT-4’s congruence with human judgments on summarization tasks (Liu et al., 2023a), thereby obviating the need for manual evaluation of CNNDM.
+Following Gao et al.
+(2023), in ASQA, we adopt the EM-recall automatic metric, which gauges the recall of correct short answers by verifying if the dataset’s provided short answers are exact substrings of the generated content, following established methodologies.
+For ELI5, we utilize the most precise automatic metric to date, claim recall, employing the TRUE (Honovich et al., 2022) natural language inference model to ascertain if the generated output encompasses the sub-claims of the reference answer.
+We present the task criteria to humans and ask them to provide a score for each answer based on the criteria.
+Participants are provided with a reference answer—not as an exclusive ground truth but as a guide—and are permitted to use search engines for additional context.
+We enlisted three annotators to evaluate 75 samples each for ASQA and ELI5, and calculate the average of them as a human score.
+Our analysis compares these human scores with those generated by task-specific metrics and GPT-4.
+The results, as detailed in our Table 4, underscore GPT-4’s closer alignment with human judgments in both ASQA and ELI5.
+As shown in Figure 12, as the task becomes more open-ended like Eli5, the human agreements become lower than the factoid dataset ASQA.
+This further evidence the assumption that the correctness of a long-form answer should be a distribution.
+ASQA ELI5
+Metric EM GPT-4 Claim GPT-4
+Corr $\uparrow$ 47.8 76.2 42.9 71.5
+MAE $\downarrow$ 43.0 12.9 52.0 9.1
+Table 4: Comparison of Human, Task-Specific metric, and GPT-4 Correctness Distributions in ASQA and ELI5 Tasks.
+Results are with %.
+‘Corr’ denotes Pearson correlation (the higher the better), ‘MAE’ denotes mean absolute error (the lower the better).
+‘EM’ denotes EM-recall, and ‘Claim’ denotes Claim-recall.
+Figure 6: Correlation for ASQA
+Figure 7: Cohen Kappa for ASQA
+Figure 8: MAE for ASQA
+Figure 9: Correlation for ELI5
+Figure 10: Cohen Kappa for ELI5
+Figure 11: MAE for ELI5
+Figure 12: Pairwise annotator agreements for ASQA and ELI5 respectively
+Appendix D Main Experiment
 We assess the calibration of variously sized models (Llama-2-13b, Llama-2-70b, Vicuna-13b, GPT-3.5-turbo) across three long-form QA datasets (ASQA, ELI5, QAMPARI) and one summarization dataset (CNNDM).
@@
 Calibration Measurement: .
-C.1 Generation
+D.1 Generation
 In the answer generation phase, we employ a 3-shot in-context learning approach for each long-form QA dataset, providing three exemplars to guide the models.
@@
 The generation of answers utilizes a top-K sampling method, setting the generation temperature for all models at 0.6 and the top-K parameter at 10.
-C.2 Correctness Evaluation
+D.2 Correctness Evaluation
 For ASQA, ELI5, CNNDM, we ask the GPT-4 to evaluate the correctness of examples.
 For ASQA and ELI5, We ask GPT-4 to evaluate the answer three times, producing a distribution capturing criteria ambiguity and model’s inherent subjectivity.
-However, for CNNDM, we only ask GPT-4 to evaluate once to save computation given that most answers correctness concentrate around 0.8 with small variance.
-The GPT-4 evaluation template can be found in Appendix E.1 and criteria for different tasks can be found in Appendix E.2.
+However, for CNNDM, we only ask GPT-4 to evaluate once to save computation given that most answers correctness concentrates around 0.8 with small variance.
+The GPT-4 evaluation template can be found in Appendix F.1 and criteria for different tasks can be found in Appendix F.2.
 For QAMPARI, which involves generating a list of entities as answers, we determine that F1 scores provide a more suitable measure of correctness than GPT-4 evaluations.
 To this end, we adopt the F1-5 metric (Gao et al., 2023), computing the F1 score based on direct matches with the correct answer list and assigning a recall of 100% for responses containing at least five accurate answers (recall-5).
-C.3 Confidence Elicitation
+D.3 Confidence Elicitation
 After getting the answer from the model, we leverage the self-evaluation or self-consistency method to derive confidence distribution from the model.
@@
 Additionally, it contains specific instructions for self-evaluation, incorporating both the question and the answer under evaluation.
-The detailed self-evaluation template and criteria are available in Appendix E.3 and Appendix E.2 respectively.
+The detailed self-evaluation template and criteria are available in Appendix F.3 and Appendix F.2 respectively.
 Self-consistency In the self-consistency approach, we generate an answer to the same question 10 times, designating the first response as the primary answer.
@@
 To assess the similarity between the two answers, we first identify factual claims in the first answer using a ClaimBuster-trained DeBERTa V2 (He et al., 2020) claim detector.
-We then verify the presence of these claims in the second answer using GPT-3.5-turbo.
+We then verify the presence of these claims in the second answer through NLI.
 The similarity score is the average presence of factual claims across answers.
@@
 For ELI5, where answers provide easily understandable explanations, we focus on the overall content.
-Here, we apply a self-consistency-naive method, asking an additional model (GPT-3.5-turbo) to assign a similarity score to each answer pair.
+Here, we apply a self-consistency-naive method, assign a similarity score to each answer pair with simple prompting.
 In CNNDM, where answers are summaries highlighting key points of an article, we gauge the similarity between two answers by evaluating the overlap of key points.
 To achieve this, we implement the self-consistency-split method.
-We dissect the first answer into individual sentences and then use GPT-3.5-turbo to determine if each sentence is present in the second answer.
+We dissect the first answer into individual sentences and then use NLI to determine if each sentence is present in the second answer.
 The similarity score is derived by averaging the presence of these segmented sentences in the comparative answer.
-Appendix D Improving Calibration
-D.1 Fine-tuning
+Appendix E Improving Calibration
+E.1 Fine-tuning
 In our study, we focus on improving model calibration on the ASQA dataset.
-We generate a self-evaluation training dataset with GPT-4, which evaluates different models responses to questions drawn from ASQA training set.
+We explore three different strategies: fine-tune the model to do self-evaluation (Input: question and model’s answer; Output: answer’s score and explanation), fine-tune the model to do generation (Input: question; Output: answer), and a hybrid of both.
+Data We generate a self-evaluation training dataset with GPT-4, which evaluates different models responses to questions drawn from ASQA training set.
+This dataset comprises inputs of questions and corresponding model answers, with outputs including a score and an explanation for each answer.
 We create the self-evaluation data in two steps.
@@
 This approach results in a comprehensive dataset with 2,000 self-evaluation examples (1,800 for training and 200 for validation), each including a question, model-generated answer, score, and justification.
-For generation data, we use 80% of ASQA’s dataset (4,353 examples) for training and the remaining 20% for validation.
+For generation data, we use 80% of training ASQA’s dataset (4,353 examples) for training and the remaining 20% for validation.
 The hybrid dataset combines the self-evaluation and generation training sets, using the self-evaluation validation set for assessment.
-Regarding training, we fine-tune Llama2-13b-chat model using LORA on this dataset.
+Training and Results Regarding training, we fine-tune Llama2-13b-chat model using LORA on this dataset.
 We maintain consistent parameters across all scenarios: a learning rate of 5e-6, five epochs, 100 warm-up steps, and a total batch size of 4 (achieved through 4 gradient accumulation steps across four GPUs, with a batch size of 1 per device).
-D.2 Hybrid Confidence Elicitation
-We integrate self-evaluation and self-consistency methods to assess their combined impact on model calibration.
-Experiments are performed on the ASQA dataset using Llama-2-13b, Llama-2-70b, and Vicuna-13b models, adjusting the self-consistency weight from 0 to 1.
-Our findings indicate that this hybrid approach improves correlation, though it does not significantly affect other metrics.
-[…]
+As Table 3 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task, as well as the limitation of LORA fine-tuning.
+Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’).
+The model becomes more confident in terms of self-consistency after fine-tuning.
+E.2 Source Documents
+We investigate the effect of additional context on LLM calibration, focusing on self-consistency confidence with Llama-2-13b-chat and GPT-3.5-turbo models on the ASQA dataset.
+We test the model’s performance when supplemented with two different types of source documents: random documents related to the question and ‘oracle’ documents directly relevant to the answers, as included in the dataset release Stelmakh et al.
+(2022).
+Findings in Table 5 reveal that Oracle documents can enhance model performance and calibration across two out of three metrics for both models, while random documents are less effective.
+These results underscore the importance of relevant contextual information in model calibration.
+Model Doc Corr ECE-M F10.8 Score
+Llama213b N 48.0-- 15.9$\uparrow$ 3.6-- 51.3--
+R 39.0 $\downarrow$ 14.8$\downarrow$ 30.6 $\uparrow$ 59.6 $\uparrow$
+O 49.5 $\uparrow$ 17.5$\uparrow$ 20.1 $\uparrow$ 65.8 $\uparrow$
+GPT-3.5turbo N 30.5-- 26.7$\uparrow$ 58.4-- 72.6$\downarrow$
+R 36.6 $\uparrow$ 26.4$\downarrow$ 68.2 $\uparrow$ 64.8 $\downarrow$
+O 27.0 $\downarrow$ 24.9$\downarrow$ 72.8 $\uparrow$ 75.8 $\uparrow$
+Table 5: How source documents affect the calibration score.
+In the document column, “N” means no documents in the input prompt, “R” means randomly selected documents relevant to the topic, and “O” means the oracle documents relevant to the answer.
+$\uparrow$ denotes that the calibration score goes up when adding documents, while $\downarrow$ means going down.
+For ECE-M it’s opposite.
+E.3 Hybrid Confidence Elicitation
+We explore whether combining self-evaluation and self-consistency can yield a more accurate confidence distribution on ASQA dataset.
+By blending confidence distributions from self-evaluation ($C^{eval}_{i}$) and self-consistency ($C^{consist}_{i}$) into a hybrid distribution $C^{hybrid}_{i}=\alpha C^{eval}_{i}+(1-\alpha)C^{consist}_{i}$, we adjust their relative contributions using $\alpha$.
+As shown in Figure 13, we observe that the correlation between confidence and correctness initially increases but then declines as $\alpha$ varies from 0 to 1.
+However, this trend doesn’t extend to other metrics like F1, indicating that while hybrid calibration elicitation may enhance calibration in terms of correlation, it may not have the same impact on other dimensions.
 […]
@@
 Figure 13: Hybrid confidence elicitation
-D.3 LLM Cascade
-Our research utilizes ASQA and QAMPARI datasets, applying self-consistency confidence in our methodology.
-We use the initial 500 samples from the evaluation data provided in Gao et al.
-(2023) as our test set and samples from 700 to 800 as the validation set.
-Initially, we assess the calibration of models Llama-2-13b, Llama-2-70b, and Llama-2-70b on the validation set, with the findings presented in Table 4.
-Notably, the Llama-2-70b model exhibits inferior calibration compared to the other models when tested on QAMPARI, leading us to exclude it from the model cascade for the test set.
-In the cascading process, we prioritize models based on their performance, ensuring that the most capable model addresses questions that multiple models confidently identify.
-Consequently, the sequence of cascaded models for ASQA is Llama-2-70b, Llama-2-13b, Vicuna-13b, GPT-3.5-turbo, and for QAMPARI, it is Llama-2-13b, Vicuna-13b, GPT-3.5-turbo.
-Then we perform our pipeline on the test sets for these tasks.
-[…]
-ASQA
-Vicuna-13b 50.3 70.7 49.2 47.4
-Llama-2-13b 51.7 76.4 21.3 47.6
-Llama-2-70b 48.0 76.0 45.2 60.4
-QAMPARI
-Vicuna-13b 48.8 75.5 27.3 13.5
-Llama-2-13b 43.0 66.2 45.7 14.1
-Llama-2-70b 3.4 56.0 20.0 8.9
-Table 4: Open-source models results on validation sets of ASQA and QAMPARI
-Appendix E Prompts
+Appendix F Prompts
 This section introduces the prompts used for our experiments.
-E.1 Correctness Evaluation Template
-Similar to contemporary work (Kim et al., 2023), our evaluation template for GPT-4 evaluation to get the ground-truth correctness distribution of an answer includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
-{task instruction} You will be given a question, a reference answer, and a student’s answer.
+F.1 Correctness Evaluation Template
+Similar to contemporary work (Kim et al., 2023), our evaluation template for GPT-4 evaluation to get the target correctness distribution of an answer includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
+Evaluation Template {task instruction} You will be given a question, a reference answer, and a student’s answer.
 Please evaluate the student’s answer based on both your knowledge and the reference answer, and provide a score from 0-5 to the student’s answer.
@@
 Question: {question} Reference answer: {reference answer} Student’s answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation.
-E.2 Criteria
+F.2 Criteria
 Below are the criteria for various tasks, with a special note that the CNNDM summarization task utilizes a distinct evaluation template.
 ASQA Criterion
-5 - Completely Correct and Highly Relevant: The answer fully addresses the question, resolves the ambiguity, and provides a well-rounded resolution.
+ASQA Criterion 5 - Completely Correct and Highly Relevant: The answer fully addresses the question, resolves the ambiguity, and provides a well-rounded resolution.
 All facts presented in the answer are accurate and relevant.
@@
 ELI5 Criterion
-5 - Perfectly Addressed, Accurate and Clarity: The answer flawlessly addresses the question with exceptional accuracy and clarity.
+ELI5 Criterion 5 - Perfectly Addressed, Accurate and Clarity: The answer flawlessly addresses the question with exceptional accuracy and clarity.
 It simplifies complex concepts effectively and does so in a way that is captivating and memorable.
@@
 It fails to address the question and does not simplify complex ideas.
-* Estimated Accuracy Assessment (0-3 Points) * 3 Points: All answers provided seem correct based on available knowledge.
+QAMPARI Criterion * Estimated Accuracy Assessment (0-3 Points) * 3 Points: All answers provided seem correct based on available knowledge.
 * 2 Points: Majority of the answers seem correct.
@@
 * Total Score (0-5 Scale) * Add the points from Estimated Accuracy and Estimated Completeness.
-You will be given one summary written for a news article.
+Summarization Evaluation Template You will be given one summary written for a news article.
 Your task is to rate the overall quality of the summary with a score from 0 to 5, where 0 is the lowest and 5 is the highest.
@@
 Task Input: Article: {article} Summary: {summary} Now please provide your score of the summary in the format of “Score: <Your score>/5” and give your explanation.
-E.3 Self-Evaluation Template
-{task instruction} You will be given a question and a student’s answer.
+F.3 Self-Evaluation Template
+Self Evaluation Template {task instruction} You will be given a question and a student’s answer.
 Please evaluate the student’s answer, and provide a score from 0-5 to the student’s answer based on the following scoring criterion: {criterion} Here are some examples.
@@
 Assess for both factual accuracy and relevance to the question.
-E.4 Self-Consistency Template
+F.4 Self-Consistency Template
 In self-consistency approaches, assessing the similarity between two answers requires the use of an additional LLM in naive and split/claim methods.
@@
 The templates for both the naive method and the split/claim method are outlined below.
-You will be presented with a question followed by two answers.
+Similarity Comparison Naive Template You will be presented with a question followed by two answers.
 Evaluate how similar these answers are, considering their amount of information provided, factual content, effectiveness in addressing the question, format and organization.
@@
 Make your answer short and concise.
-You will receive a sentence and a response; please ascertain if a similar statement is present in the response.
+Similarity Comparison Split Template You will receive a sentence and a response; please ascertain if a similar statement is present in the response.
 Sentence: {sentence} Response: {response} Verify if the response contains a statement resembling the target sentence.
 Answer yes or no.
-Appendix F Qualitative Analysis
-We showcase some successful and failed calibration examples from LLama-2-13b on Qampari and ASQA with self-consistency method.
+Appendix G Qualitative Analysis
+We showcase successful and failed calibration examples on Qampari and ASQA with self-consistency method.
 type content
 QAMPARI Successful Example 1
-Question
-Who directed a film that had P. Balachandran as a screenwriter?
-Answer
-Kamal, P. Balachandran, T. K. Rajeev Kumar, V. K. Prakash, Rajeev Ravi.
-Model Answer
-Adu, Ammayum Muyalum, Bhoothakkannadi, Chattambi, Kakkakku Pinnaka…
-Other Answer 1
-Agraharathile Kazhuthai, Kaliyugam, Kodungallur, Kodiyettam …
-Other Answer 2
-Kodiyettam, Kuttyedam, Oru CBI Diary Kurippu…
-Other Answer 3
-Aravindan, K.G. George, John Abraham, Shyam Benegal …
-Self-consistency scores
-[0, 0, 0, …] (%)
-Correctness
-Score: $0\%$ Distribution: $\{0:1.0\}$
-Confidence score
-Score: $0\%$ Distribution: $\{0:1.0\}$
+Question Who directed a film that had P. Balachandran as a screenwriter?
+Answer Kamal, P. Balachandran, T. K. Rajeev Kumar, V. K. Prakash, Rajeev Ravi.
+Model Answer Adu, Ammayum Muyalum, Bhoothakkannadi, Chattambi, Kakkakku Pinnaka…
+Other Answer 1 Agraharathile Kazhuthai, Kaliyugam, Kodungallur, Kodiyettam …
+Other Answer 2 Kodiyettam, Kuttyedam, Oru CBI Diary Kurippu…
+Other Answer 3 Aravindan, K.G. George, John Abraham, Shyam Benegal …
+Self-consistency scores [0, 0, 0, …] (%)
+Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
+Confidence score Score: $0\%$ Distribution: $\{0:1.0\}$
 QAMPARI Successful Example 2
-Question
-Which cities or municipalities are located in the province of Basilan?
-Answer
-Akbar, Al-Barka, Hadji Muhtamad, Isabela City, Lamitan, Lantawan, Maluso, Sumisip, Tabuan-Lasa, Tipo-Tipo, Tuburan, Ungkaya Pukan.
-Model Answer
-Al-Barka, Basilan, Isabela, Lamitan, Tipo-Tipo, Maluso, Sumisip.
-Other Answer 1
-Basilan, Isabela, Lamitan, Lantawan, Tipo-Tipo, Sumisip.
-Other Answer 2
-Basilan, Tipo-Tipo, Isabela City, Maluso, Sumisip, Tuburan.
-Other Answer 3
-Basilan, Isabela, Tipo-Tipo, Lamitan, Sumisip, Al-Barka, Maluso, Tuburan.
-Self-consistency scores
-[71.4, 57.1, 85.7, …](%)
-Correctness
-Score: $83.3\%$ Distribution: $\{0.8:1.0\}$
-Confidence
-Score: $76.2\%$ Distribution: $\{0.6:0.11,0.8:0.78,1.0:0.11\}$
+Question Which cities or municipalities are located in the province of Basilan?
+Answer Akbar, Al-Barka, Hadji Muhtamad, Isabela City, Lamitan, Lantawan, Maluso, Sumisip, Tabuan-Lasa, Tipo-Tipo, Tuburan, Ungkaya Pukan.
+Model Answer Al-Barka, Basilan, Isabela, Lamitan, Tipo-Tipo, Maluso, Sumisip.
+Other Answer 1 Basilan, Isabela, Lamitan, Lantawan, Tipo-Tipo, Sumisip.
+Other Answer 2 Basilan, Tipo-Tipo, Isabela City, Maluso, Sumisip, Tuburan.
+Other Answer 3 Basilan, Isabela, Tipo-Tipo, Lamitan, Sumisip, Al-Barka, Maluso, Tuburan.
+Self-consistency scores [71.4, 57.1, 85.7, …](%)
+Correctness Score: $83.3\%$ Distribution: $\{0.8:1.0\}$
+Confidence Score: $76.2\%$ Distribution: $\{0.6:0.11,0.8:0.78,1.0:0.11\}$
 QAMPARI Failed Example 1 (Under Confident)
-Question
-What city or municipality located in the Philipines are in the Tarlac province?
-Answer
-Anao, Bamban, Camiling, Capas, Concepcion, Gerona, La Paz, Mayantoc, Moncada, Paniqui, Pura, Ramos, San Clemente, San Jose, San Manuel, Santa Ignacia, Tarlac City, Victoria.
-Model Answer
-Capas, Concepcion, Gerona, La Paz, Mayantoc, Paniqui, San Jose, Santa Ignacia, Tarlac City, Victoria.
-Other Answer 1
-Capas, Concepcion, Camiling, Tarlac City, Paniqui, San Jose, Santa Lucia, Victoria.
-Other Answer 2
-Capas, Tarlac, Concepcion, Tarlac, Gerona, Tarlac, La Paz, Tarlac, Paniqui, Tarlac, San Jose, Tarlac, Tarlac City.
-Other Answer 3
-Capas, Concepcion, Camiling, Tarlac City, San Jose.
-Self-consistency scores
-[60, 60, 30, …](%)
-Correctness
-Score: $100\%$ Distribution: $\{1:1.0\}$
-Confidence
-Score: $51.1\%$ Distribution: $\{0.4:0.56,0.6:0.33,0.8:0.11\}$
+Question What city or municipality located in the Philipines are in the Tarlac province?
+Answer Anao, Bamban, Camiling, Capas, Concepcion, Gerona, La Paz, Mayantoc, Moncada, Paniqui, Pura, Ramos, San Clemente, San Jose, San Manuel, Santa Ignacia, Tarlac City, Victoria.
+Model Answer Capas, Concepcion, Gerona, La Paz, Mayantoc, Paniqui, San Jose, Santa Ignacia, Tarlac City, Victoria.
+Other Answer 1 Capas, Concepcion, Camiling, Tarlac City, Paniqui, San Jose, Santa Lucia, Victoria.
+Other Answer 2 Capas, Tarlac, Concepcion, Tarlac, Gerona, Tarlac, La Paz, Tarlac, Paniqui, Tarlac, San Jose, Tarlac, Tarlac City.
+Other Answer 3 Capas, Concepcion, Camiling, Tarlac City, San Jose.
+Self-consistency scores [60, 60, 30, …](%)
+Correctness Score: $100\%$ Distribution: $\{1:1.0\}$
+Confidence Score: $51.1\%$ Distribution: $\{0.4:0.56,0.6:0.33,0.8:0.11\}$
 QAMPARI Failed Example 1 (Over Confident)
-Question
-What are some notables mailing list softwares ?
-Answer
-Dada Mail, Discourse, GNU Mailman, GroupServer, LISTSERV, Majordomo, Organic Groups Mailing List, Sympa.
-Model Answer
-Mailchimp, Constant Contact, Campaign Monitor, AWeber, GetResponse, MailPoet, Sendinblue, MailerLite, Klaviyo, ConvertKit.
-Other Answer 1
-Mailchimp, AWeber, GetResponse, Campaign Monitor, Constant Contact, Klaviyo, ConvertKit, MailPoet, Mailster, Sendinblue, MailerLite.
-Other Answer 2
-Mailchimp, Constant Contact, AWeber, Campaign Monitor, GetResponse, Klaviyo, MailPoet, Mailster, Sendinblue, MailerLite, ConvertKit, ActiveCampaign, and Constant Contact.
-Other Answer 3
-Mailchimp, AWeber, GetResponse, Campaign Monitor, MailerLite, Constant Contact, Klaviyo, ConvertKit, Sendinblue, and Emma.
-Self-consistency scores
-[100, 100, 90, …](%)
-Correctness
-Score: $0\%$ Distribution: $\{0:1.0\}$
-Confidence score
-Score: $97.8\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
+Question What are some notables mailing list softwares ?
+Answer Dada Mail, Discourse, GNU Mailman, GroupServer, LISTSERV, Majordomo, Organic Groups Mailing List, Sympa.
+Model Answer Mailchimp, Constant Contact, Campaign Monitor, AWeber, GetResponse, MailPoet, Sendinblue, MailerLite, Klaviyo, ConvertKit.
+Other Answer 1 Mailchimp, AWeber, GetResponse, Campaign Monitor, Constant Contact, Klaviyo, ConvertKit, MailPoet, Mailster, Sendinblue, MailerLite.
+Other Answer 2 Mailchimp, Constant Contact, AWeber, Campaign Monitor, GetResponse, Klaviyo, MailPoet, Mailster, Sendinblue, MailerLite, ConvertKit, ActiveCampaign, and Constant Contact.
+Other Answer 3 Mailchimp, AWeber, GetResponse, Campaign Monitor, MailerLite, Constant Contact, Klaviyo, ConvertKit, Sendinblue, and Emma.
+Self-consistency scores [100, 100, 90, …](%)
+Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
+Confidence score Score: $97.8\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
 Table 6: Examples for QAMPARI
@@
 ASQA Successful Example 1
-Question
-What are the three types of enzymes in digestion?
-Answer
-Digestive enzymes are a group of enzymes that break down polymeric… Lipases split fatty acids… Proteases and peptidases split proteins … Amylases split carbohydrates … Nucleases split nucleic acids into nucleotides…
-Model Answer
-There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats.
+Question What are the three types of enzymes in digestion?
+Answer Digestive enzymes are a group of enzymes that break down polymeric… Lipases split fatty acids… Proteases and peptidases split proteins … Amylases split carbohydrates … Nucleases split nucleic acids into nucleotides…
+Model Answer There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats.
 Proteases are produced by the stomach and pancreas, while amylases are … These enzymes work in concert …
-Other Answer 1
-Digestive enzymes are proteins that help … There are three main types of digestive enzymes: proteases, which break down proteins into amino acids; amylases, which break down carbohydrates into simple sugars; and lipases, which break down fats into glycerol and fatty acids.
+Other Answer 1 Digestive enzymes are proteins that help … There are three main types of digestive enzymes: proteases, which break down proteins into amino acids; amylases, which break down carbohydrates into simple sugars; and lipases, which break down fats into glycerol and fatty acids.
 These enzymes are produced by different organs and …
-Other Answer 2
-There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats.
+Other Answer 2 There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats.
 These enzymes are produced by the salivary glands, stomach, and pancreas, and work in various parts of the digestive system…
-Self-consistency scores
-[100, 100, 0, …] (%)
-Correctness
-Score: $100\%$ Distribution: $\{1.0:1.0\}$
-Confidence score
-Score: $95.6\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
+Self-consistency scores [100, 100, 0, …] (%)
+Correctness Score: $100\%$ Distribution: $\{1.0:1.0\}$
+Confidence score Score: $95.6\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
 ASQA Successful Example 2
-Question
-When did ben rectors new album come out?
-Answer
-Benjamin Evans Rector is an American singer… He has released seven studio albums, including Magic on June 22, 2018, which peaked at No.
+Question When did ben rectors new album come out?
+Answer Benjamin Evans Rector is an American singer… He has released seven studio albums, including Magic on June 22, 2018, which peaked at No.
 1 ..
 Brand New which released on August 28, 2015… Older albums such as Something Like This, released on September 13, 2011 and The Walking in Between, came out August 20, 2013.
-Model Answer
-Ben Rector’s new album, "Magic", was released on March 25, 2022.
-Other Answer 1
-Ben Rector’s new album, "Brand New", was released on March 10th, 2022.
-Other Answer 2
-Ben Rector, an American singer-songwriter, released his new album "Magic" on October 20, 2022.
-Self-consistency scores
-[0, 0, …](%)
-Correctness
-Score: $0\%$ Distribution: $\{0:1.0\}$
-Confidence
-Score: $0\%$ Distribution: $\{0:1.0\}$
+Model Answer Ben Rector’s new album, "Magic", was released on March 25, 2022.
+Other Answer 1 Ben Rector’s new album, "Brand New", was released on March 10th, 2022.
+Other Answer 2 Ben Rector, an American singer-songwriter, released his new album "Magic" on October 20…
+Self-consistency scores [0, 0, …](%)
+Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
+Confidence Score: $0\%$ Distribution: $\{0:1.0\}$
 ASQA Failed Example 1 (Under Confident)
-Question
-Who plays charlie on it’s always sunny?
-Answer
-Charlie Day plays the fictional character Charlie Kelly… Day is an American actor, screenwriter…
-Model Answer
-Charlie Day plays Charlie on It’s Always Sunny in Philadelphia. He is also known for his role as Dale Arbus … Day is a member of the improvisational comedy …
-Other Answer 1
-Dennis is played by Glenn Howerton.
-Other Answer 2
-Charlie on It’s Always Sunny in Philadelphia is played by Charlie Day.
-Self-consistency scores
-[0, 33.3, …](%)
-Correctness
-Score: $80\%$ Distribution: $\{0.8:1.0\}$
-Confidence
-Score: $31.1\%$ Distribution: $\{0:0.22,0.4:0.78\}$
+Question Who plays charlie on it’s always sunny?
+Answer Charlie Day plays the fictional character Charlie Kelly… Day is an American actor, screenwriter…
+Model Answer Charlie Day plays Charlie on It’s Always Sunny in Philadelphia. He is also known for his role as Dale Arbus … Day is a member of the improvisational comedy …
+Other Answer 1 Dennis is played by Glenn Howerton.
+Other Answer 2 Charlie on It’s Always Sunny in Philadelphia is played by Charlie Day.
+Self-consistency scores [0, 33.3, …](%)
+Correctness Score: $80\%$ Distribution: $\{0.8:1.0\}$
+Confidence Score: $31.1\%$ Distribution: $\{0:0.22,0.4:0.78\}$
 ASQA Failed Example 2 (Over Confident)
-Question
-Who sings wake me up when it’s all over?
-Answer
-“Wake Me Up” is a song by Swedish DJ and record producer Avicii.
+Question Who sings wake me up when it’s all over?
+Answer “Wake Me Up” is a song by Swedish DJ and record producer Avicii.
 Aloe Blacc, which is the stage name for Egbert Nathaniel Dawkins III provides vocals for the song while Mike Einziger provides acoustic guitar.
-“Wake Me When It’s Over” is a song written and recorded by country singer Willie Nelson in 1962.
-Model Answer
-“Wake Me Up When September Ends” is a song written by Billy
-Other Answer 1
-“Wake Me Up When September Ends” is a song by American rock band Green Day, written by lead vocalist Billie Joe Armstrong…
-Other Answer 2
-“Wake Me Up When September Ends” is a song by Green Day, written by the band’s lead vocalist and guitarist Billie Joe Armstrong…
-Self-consistency scores
-[100, 100, …](%)
-Correctness
-Score: $0\%$ Distribution: $\{0:1.0\}$
-Confidence score
-Score: $77.8\%$ Distribution: $\{0:0.22,1.0:0.78\}$
+“Wake Me When It’s Over” is a song written and recorded by Willie…
+Model Answer “Wake Me Up When September Ends” is a song written by Billy
+Other Answer 1 “Wake Me Up When September Ends” is a song by American rock band Green Day, written by lead vocalist Billie Joe Armstrong…
+Other Answer 2 “Wake Me Up When September Ends” is a song by Green Day, written by the band’s lead vocalist and guitarist Billie Joe Armstrong…
+Self-consistency scores [100, 100, …](%)
+Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
+Confidence score Score: $77.8\%$ Distribution: $\{0:0.22,1.0:0.78\}$
 Table 7: Examples for ASQA